### v2

# Towards a Cascaded LLM Framework for Cost-effective Human-AI Decision-Making

Abstract
Effective human-AI decision-making balances three key factors: the correctness of predictions, the cost of knowledge and reasoning complexity, and the confidence about whether to abstain automated answers or involve human experts. In this work, we present a cascaded LLM decision framework that adaptively delegates tasks across multiple tiers of expertise – a base model for initial candidate answers, a more capable and knowledgeable (but costlier) large model, and a human expert for when the model cascade abstains. Our method proceeds in two stages. First, a deferral policy determines whether to accept the base model’s answer or regenerate it with the large model based on the confidence score. Second, an abstention policy decides whether the cascade model response is sufficiently certain or requires human intervention. Moreover, we incorporate an online learning mechanism in the framework that can leverage human feedback to improve decision quality over time. We demonstrate this approach to general question-answering (ARC-Easy and ARC-Challenge) and medical question-answering (MedQA and MedMCQA). Our results show that our cascaded strategy outperforms in most cases single-model baselines in accuracy while reducing cost and providing a principled way to handle abstentions. [^1] We provide the code for our experiments at https://github.com/fanconic/cascaded-llms
Machine Learning, ICML
1 Introduction
Data-driven decision support has gained increasing traction in high-stakes fields such as healthcare (Jin et al., 2024; Fan et al., 2024; Li et al., 2024), finance (Li et al., 2023a; Zhao et al., 2024), and education (Xu et al., 2024). For example, in the medical context, large language models (LLMs) can facilitate accurate diagnoses and treatment recommendations that encode vast knowledge (Kim et al., 2024). However, high accuracy in such complex settings often requires substantial computational resources or multiple reasoning steps. Additionally, LLMs may hallucinate or generate incorrect outputs with severe consequences. Effective human-AI collaboration should balance correctness, cost, and abstention, ensuring AI-driven assistance integrates seamlessly with expert oversight.
Challenge:
Efficient allocation of computational and human resources is essential, as larger models incur significantly higher costs, especially with advanced reasoning methods, like searching the solution space (OpenAI, 2025; Anthropic, 2025). Decision-making tasks differ in complexity. Smaller models can handle routine cases, while complex queries require more advanced reasoning. Choosing between deploying a small model, a more sophisticated model, or referring to a human expert is non-trivial, particularly under resource constraints. A static approach — always using a large model or defaulting to small ones — fails to optimise cost-benefit trade-offs.
Figure 1: Cascaded LLM Human-AI Decision-Making Framework. Given a decision-making problem, the system (1) generates an initial response with a base model, (2) verifies correctness probability, (2.5) defers to a larger model if needed, (3) assesses response uncertainty, and (3.5) abstains to a human expert if necessary. If feedback is available, deferral and abstention modules are adjusted over time. For this system to work efficiently, the modules should uphold three desiderata: D1 the deferral policy regenerates responses only when necessary, D2 the abstention policy escalates to humans only when uncertainty is high, D3 the system continuously improves with feedback.
Our Solution:
We propose a cascaded multi-tier framework integrating (1) a cost-effective base model, (2) an advanced but resource-intensive large model, and (3) a human expert in case of abstention. This system prioritises efficiency, deferring only when likely and necessary. A deferral policy determines if the large model is required to regenerate the answer, while an abstention policy forwards the query to experts when uncertainty is high. Online learning continuously refines these policies based on feedback. Figure 1 illustrates how decisions are handled across datasets of varying difficulty.
Framework Desiderata:
Three desiderata must be upheld for our proposed framework to function correctly:
D1 : Reduce Unnecessary Regenerations The deferral policy should escalate to the large model only if the base model’s response is likely incorrect, avoiding unnecessary computational costs.
D2 : Abstain and Escalate to Humans when Uncertain The abstention policy should refer to human experts only when the cascade system’s uncertainty is too high, reducing unnecessary human intervention.
D3 : Dynamic Improvement based on Feedback The framework should continually refine its policies based on feedback, ensuring adaptive and risk-minimised decision-making.
Contributions:
We make the following key contributions in this paper:
  • Cascaded LLM human-AI decision system: A framework integrating LLMs of varying complexity with human experts for accuracy, efficiency, and uncertainty management.
  • Investigation into Cost-Efficient Verification Methods: We analyse different LLM verification methods to determine the most suited for cost-efficient deferral.
  • Online learning framework: An online learning approach that continuously improves deferral and abstention policies when feedback is available, enhancing decision-making adaptively over time.
2 Related Work
Multi-LLM Answer Generation. Several studies have explored collaborative frameworks that leverage multiple LLMs of varying capacities to enhance both performance and cost-efficiency beyond the capabilities of a single model (Chen et al., 2023; Ding et al., 2024; Aggarwal et al., 2024). (Chen et al., 2023) proposed cost-effective strategies such as prompt structuring, model approximation, and cascaded LLM frameworks. Similarly, (Ding et al., 2024) introduced an intelligent routing mechanism that dynamically assigns prompts to the most appropriate model. (Aggarwal et al., 2024) developed a black-box LLM framework for cost-efficient response generation, formalised as a Partially Observable Markov Decision Process (POMDP), requiring minimal training data. (Zhu et al., 2023a) proposed a multiplexer-based approach that balances queries between a small and a large LLM, employing a trained BERT classifier to determine when the smaller model suffices. (Šakota et al., 2024) introduced a meta-model-driven selection framework that requires pre-training for optimal query distribution. In a parallel line of research, speculative decoding (Leviathan et al., 2023), employs a lightweight model to generate multiple tokens, which a larger model subsequently verifies.
In contrast to prior research, we propose a multi-tier framework for human-AI collaboration. Rather than relying solely on automation, our approach integrates human intervention when model uncertainty is too high, addressing a gap in previous multi-tier frameworks. Compared to speculative decoding research, our work prioritises the factual correctness of complete responses rather than token-wise distributions, enabling more robust decision-making rather than just fluent text generation. (Zellinger et al., 2025) conducts a concurrent line of research that is closest to our work on cascaded LLMs, as well as in their previous works (Zellinger and Thomson, 2024, 2025). They focus on probabilistic modelling of cascading LLMs and their deferral and abstention mechanisms.
LLM Answer Verification and Uncertainty Quantification. Ensuring the reliability of LLM-generated responses requires adequate verification and uncertainty quantification mechanisms. Several studies have explored self-verification strategies (Weng et al., 2023; Jiang et al., 2024; Pan et al., 2024), often leveraging the LLM’s internal knowledge (Dhuliawala et al., 2023). Alternative approaches employ external knowledge sources for verification (Pan et al., 2024; Gao et al., 2023; Peng et al., 2023). (Aggarwal et al., 2024) introduced verification techniques based on available contextual information, predominantly involving multiple LLM queries to validate response accuracy. Another research direction quantifies factual correctness uncertainty (Mahaut et al., 2024). (Kadavath et al., 2022) conducted a detailed analysis of how LLMs express uncertainty through surrogate token probabilities, demonstrating their effectiveness in calibration. (Azaria and Mitchell, 2023) explored internal LLM states, training classifiers to quantify uncertainty, while methods such as semantic uncertainty estimation (Kuhn et al., 2023) enhance robustness by analysing variations in semantically equivalent token sequences.
Our approach relies on surrogate token probability (Kadavath et al., 2022) as a core verification component. However, we extend this methodology by integrating a hierarchical escalation mechanism that dynamically transitions between models and human experts based on verification results.
Selective prediction. Selective prediction enables models to abstain from uncertain queries (El-Yaniv and Wiener, 2010), a crucial feature in risk-sensitive settings where errors are costly. The idea dates back to Chow’s work on optical character recognition (Chow, 1957, 1970), and has since been shown to improve deep learning performance (Geifman and El-Yaniv, 2017). In NLP, abstention has been introduced through confidence-based thresholds (Xin et al., 2021; Yoshikawa and Okazaki, 2023), with recent work on uncertainty quantification for large language models advancing this line of research (Manakul et al., 2023; Farquhar et al., 2024; Lin et al., 2024).
LLMs in Online Learning. Traditional LLM research predominantly evaluates language models on static datasets. However, our work aligns with online learning paradigms, wherein policies are continuously refined in response to streaming data (Cortes et al., 2018; Ye et al., 2024). Our methodology is inspired by (Jarrett et al., 2022), who introduced an online decision mediation framework mediating between suboptimal human decisions and an expert oracle. A similar research with the online learning approach is conducted by (Zhu et al., 2023a), which extended their multiplexer mechanism to an online setting.
3 Background
3.1 Cascaded Decision System
We consider a two-tiered cascaded LLM decision system for question answering under resource constraints, denoted by $C=\mathcal{M}_{\text{base}}\rightarrow\mathcal{M}_{\text{large}}$, following the notation of Zellinger et al. (2025). Let $x\in\mathcal{X}$ be a problem statement or prompt, and let $y\in\mathcal{Y}$ denote a system-generated response. For every input $x$, the models return a confidence score $\Phi_{i}(x)\in[0,1]$ and an uncertainty score $\Xi_{i}(x)\in[0,\infty)$, where $i\in\{\text{base},\text{large}\}$. The decision to predict using the base model $\mathcal{M}_{\text{base}}$ or to defer to the larger model $\mathcal{M}_{\text{large}}$ is based on whether the confidence exceeds a deferral threshold, i.e., $\Phi(x)>\phi_{\text{base}}$. Thus, a prediction is only made if the base model is sufficiently confident. In contrast, abstention is governed by predictive uncertainty: if this exceeds a threshold, $\Xi_{i}(x)>\xi_{i}$, the system abstains and forwards the query to a human expert.
We formally define the cascaded decision system as:
$\tiny C(x)=\begin{cases}\mathcal{M}_{\text{base}}(x)&\text{if }\Phi_{\text{base}}(x)>\phi_{\text{base}}\land\Xi_{\text{base}}(x)<\xi_{\text{base}}\\ \mathcal{M}_{\text{large}}(x)&\text{if }\Phi_{\text{base}}(x)\leq\phi_{\text{base}}\land\Xi_{\text{base}}(x)\leq\xi_{\text{base}}\land\Xi_{\text{large}}(x)\leq\xi_{\text{large}}\\ \varnothing&\text{if }\Xi_{\text{base}}(x)\geq\xi_{\text{base}}\,\lor\,\Xi_{\text{base}}(x)\geq\xi_{\text{base}}\end{cases}$ (1)
The decision flow of this cascade is also illustrated in Figure 2. While we focus on a two-model system here, the framework naturally generalises to cascades involving multiple LLMs of varying sizes.
The objective is to generate accurate responses while accounting for the computational costs of the models and abstaining when the system is too uncertain. As described in Zellinger and Thomson (2024), this constitutes a multi-objective optimisation problem over three dimensions: error, cost, and abstention. Formally, we minimise the system risk:
$\mathcal{R}(C)=\mathbb{P}(\text{error}\land\neg\text{abstention})+\lambda_{c}\mathbb{E}[\text{Cost}]+\lambda_{a}\mathbb{P}(\text{abstention})$ (2)
Here, $\mathbb{P}(\text{error}\land\neg\text{abstention})$ denotes the probability of the system making an error when it does not abstain, $\mathbb{E}[\text{Cost}]$ is the expected computational cost incurred, and $\mathbb{P}(\text{abstention})$ is the probability of the system abstaining and deferring to a human expert. The terms $\lambda_{c}$ and $\lambda_{a}$ weight the cost and abstention penalties, respectively. We explain the system risk in more detail in Section 4.2.
Assumptions.
A1
The base model is cost-efficient but less accurate, whereas the large model is more capable but computationally expensive. A2 Generating responses incurs significantly higher cost than processing inputs, especially in settings that require Chain-of-Thought (CoT) prompting (Wei et al., 2022) or advanced test-time reasoning (Xie et al., 2024). A3 Each response is assumed to be either correct or incorrect, with no ambiguity.
Figure 2: Decision flow of the two-tiered cascaded LLM system. A query is first evaluated by the base model. The base model response is returned if the confidence is above a deferral threshold and uncertainty is below an abstention threshold. Otherwise, the query is passed to the large model, producing a response or abstaining if uncertainty remains high. In such cases, the query is deferred to a human expert. Online feedback can be incorporated at each stage to improve system performance over time.
3.2 Cost Calculation
To estimate the computational cost incurred by a language model when generating a response, we define a cost function that scales linearly with the number of parameters and the number of tokens processed. Let $s\in\mathbb{R}_{>0}$ denote the size of the model in billions of parameters, extracted from the model identifier (e.g., “Llama-8B” $\Rightarrow$ $s=8$). Let $t_{\text{in}}$ and $t_{\text{out}}$ denote the number of input and output tokens, respectively, and let $\rho>0$ denote the output-to-input token price ratio, which accounts for the higher cost of generating tokens. The total cost is given by:
$\text{Cost}(s,t_{\text{in}},t_{\text{out}},\rho)=s\cdot\left(t_{\text{in}}+\rho\cdot t_{\text{out}}\right)$ (3)
This abstraction provides a simple yet effective means of comparing models of different sizes and outputs under a unified cost metric, without relying on specific infrastructure-dependent pricing.
4 Methods
4.1 Calibrated Confidence and Uncertainty Estimation
Effective deferral and abstention decisions in a cascaded system critically depend on accurately quantifying model confidence $\Phi(x)$ and uncertainty $\Xi(x)$ for each input $x$. Overconfident or miscalibrated predictions can lead to errors, while excessive uncertainty may result in unnecessary escalations. Therefore, the first part of our method focuses on analysing a range of techniques to estimate these quantities in a reliable and cost-efficient manner. To this end, we evaluate four complementary methods that approximate the probability that a response is correct.
(1) Self-Verification.
Given an input $x$ and a model response $y_{\text{base}}\sim\mathcal{M}_{i}(x)$, we prompt the same model to quantify how likely the response is correct by generating a new response (Li et al., 2023b). The model returns raw confidence score by outputting either a scalar value token in response to a verification prompt (see Appendix B.1). The outputted probability serves as an uncalibrated estimate of correctness.
(2) Consistent Self-Verification.
We repeat the self-verification process $n$ times under stochastic sampling (e.g., with temperature), and aggregate the resulting probabilities. The empirical mean forms the uncalibrated confidence score. This approach is inspired by self-consistency as in (Aggarwal et al., 2024).
(3) Surrogate Token Probability.
We adopt the approach of Kadavath et al. (2022), where the model $\mathcal{M}_{i}$ is asked to verify whether a generated response $y$ is correct, and we extract the next-token probability over the discrete label set YES/NO. Specifically:
$p_{i}(x)=\frac{\mathcal{M}_{i}(\texttt{YES}\mid x,y)}{\mathcal{M}_{i}(\texttt{YES}\mid x,y)+\mathcal{M}_{i}(\texttt{NO}\mid x,y)},$ (4)
(4) Monte-Carlo Surrogate Token Probability.
To obtain better confidence estimates, we apply Monte Carlo Dropout (Gal and Ghahramani, ) at test time when computing the surrogate token probability. For each of $n$ stochastic forward passes, we sample an estimate $\hat{p}^{(t)}_{i}(x)$, and the empirical mean forms the uncalibrated confidence score:
${p}_{i}(x)=\frac{1}{T}\sum_{t=1}^{n}\hat{p}^{(t)}_{i}(x)$ (5)
Model Evaluation by Larger Models.
For each of the above methods, the evaluating model $\mathcal{M}_{i}$ can either be the same model that generated the original response, or a larger model in the cascade, if available. While self-evaluation is cheap and self-contained, verifying a small model’s output using a larger model is still substantially cheaper than generating a new response from scratch—particularly when generation involves long-form reasoning, as per Assumption  A2. Additionally, larger models tend to be better calibrated and may yield more reliable verification, improving downstream deferral and abstention decisions (Zhu et al., 2023b; Chhikara, 2025).
Bayesian Calibration.
To ensure that the extracted confidence scores are comparable across models and consistent with empirical correctness, we fit a Bayesian logistic regression model on a small calibration set of 100 samples. This is a Bayesian version of Platt scaling (Platt, 2000), and we assume a Normal distribution as prior. We follow (Zellinger and Thomson, 2024)’s approach and apply a non-linear transformation on the raw confidence score before inputting it into the Bayesian model, to spread out the clusters of overconfident probabilities.
$p_{tr}(p_{i})=\begin{cases}\log(\frac{1}{1-p_{i}})&\text{if }p_{i}\geq 0.5\\ \log(2)-\log(\frac{1}{p_{i}})&\text{if }p_{i}<0.5\end{cases}$ (6)
Subsequently, the Bayesian Logistic Regression outputs a posterior distribution over correctness. The mean of the posterior predictive distribution defines the calibrated confidence $\Phi(x)$, while we use standard deviation as a model-based uncertainty estimate $\Xi(x)$, as in (Fanconi et al., 2023).
4.2 Online Improvement
To enable online learning D3, we parameterise the deferral and abstention thresholds and optimise them online. Given a dataset $\mathcal{D}^{(t)}$ at time $t\in\mathbb{N}$ with previous problem statements and ground truth labels, we update the thresholds using stochastic gradient descent. While the system is deployed, we assume that we will receive a ground truth response ($y^{*}$) at the end of every decision if the system abstains. Thus, our dataset continually increases $\mathcal{D}^{(t)}=\mathcal{D}^{(t-1)}\cup\{x,y^{*}\}$ every time the cascade abstains.
Our objective function is the system risk $\mathcal{R}(C)$ (Equation 2). We expand this risk into the concrete, differentiable losses. Throughout, let
$\Phi_{i}(x)\in[0,1],\qquad\Xi_{i}(x)\in[0,1],\qquad i\in\{\text{base},\text{large}\}$
denote the calibrated probability of correctness (posterior predictive) and a uncertainty score (i.e. posterior predictive standard deviation) returned by model $i$ for an input $x$. The optimisation variables are
$\phi_{\text{base}},\;\xi_{\text{base}},\;\xi_{\text{large}}\in(0,1),$
For numerical stability we treat their raw, unconstrained versions $\phi_{\text{base}}^{\text{raw}},\xi_{\text{base}}^{\text{raw}},\xi_{\text{large}}^{\text{raw}}\in\mathbb{R}$ as the true optimisation parameters and map them to $(0,1)$ with a sigmoid function:
$\phi_{\text{base}}=\sigma\!\bigl{(}\phi_{\text{base}}^{\text{raw}}\bigr{)},\quad\xi_{\text{base}}=\sigma\!\bigl{(}\tau_{\text{base}}^{\text{raw}}\bigr{)},\quad\xi_{\text{large}}=\sigma\!\bigl{(}\tau_{\text{large}}^{\text{raw}}\bigr{)}.$
To keep the loss fully differentiable, we replace every Boolean test with a soft logistic step, where $k$ determines the steepness
$\mathbf{1}\{z>0\}\;\longrightarrow\;g_{k}(z)=\sigma(k\,z).$
With this convention the three mutually exclusive masks at the base stage are
$p_{\text{abst1}}(x)$ $=g_{k}\!\bigl{(}\Xi_{\text{base}}(x)-\xi_{\text{base}}\bigr{)},$ (7)
$m_{\text{pred1}}(x)$ $=\bigl{(}1-m_{\text{abst1}}\bigr{)}\cdot g_{k}\!\bigl{(}\Phi_{\text{base}}(x)-\phi_{\text{base}}\bigr{)},$ (8)
$m_{\text{defer1}}(x)$ $=\bigl{(}1-m_{\text{abst1}}\bigr{)}\cdot g_{k}\!\bigl{(}\phi_{\text{base}}-\Phi_{\text{base}}(x)\bigr{)},$ (9)
and the masks at the large stage are
$p_{\text{abst2}}(x)$ $=m_{\text{defer1}}(x)\cdot g_{k}\!\bigl{(}\Xi_{\text{large}}(x)-\xi_{\text{large}}\bigr{)},$ (10)
$m_{\text{pred2}}(x)$ $=m_{\text{defer1}}(x)\cdot\bigl{(}1-g_{k}(\Xi_{\text{large}}(x)-\xi_{\text{large}})\bigr{)}.$ (11)
Probability of abstention. The cascade abstains in two mutually exclusive ways, so
$\mathbb{P}(\text{abstention})\;=\;p_{\text{abst1}}\;+{p_{\text{abst2}}}.$
Expected correctness. Only the prediction masks contribute a non-zero probability of correctness; we weight each by the calibrated confidence:
$\mathbb{E}[\text{Correct}]\;=\;\mathbb{E}\!\bigl{[}m_{\text{pred1}}\cdot\Phi_{\text{base}}\bigr{]}\;+\;\mathbb{E}\!\bigl{[}m_{\text{pred2}}\cdot\Phi_{\text{large}}\bigr{]}.$
Expected cost. Let $c_{1}$ be the costs from the base model, which consist of the generation cost and the verification cost (either by itself or by a larger model). Furthermore, $c_{2}$ is the generation cost and the verification cost caused by the large model. The first term is incurred on every query; the second is incurred only if we defer:
$\mathbb{E}[\text{Cost}]\;=\;c_{1}\;+\;\mathbb{E}[m_{\text{defer1}}]\cdot c_{2}.$
System-risk objective. Substituting the three expectations above into Eq. (2) produces the differentiable loss that is back-propagated during threshold optimisation in online learning:
$\mathcal{R}(C)$ $=1-\mathbb{E}[\text{Correct}]\;+\;\lambda_{c}\,\mathbb{E}[\text{Cost}]\;+\;\lambda_{a}\,\bigl{(}p_{\text{abst1}}+p_{\text{abst2}}\bigr{)}.$ (12)
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
Figure 3: Cost-Accuracy Trade-off for Calibrated Verification Methods (Llama 1B–8B). Accuracy versus cost per sample is shown for the cascaded model using various verification methods. Performance above the linear interpolation line between base and large model baselines indicates a positive cost-benefit. Error bars represent standard error.
5 Experiments
In this section, we empirically assess whether the desiderata D1, D2, and D3, introduced in Section 1, are satisfied. For D1 and D2, we analyse in Section 5.1 the performance of various confidence estimation techniques with respect to calibration and cost-efficiency. Subsequently, in Section 5.2, we investigate whether the system improves through online learning.
General Setup.
[…]
To evaluate the generalisability of our framework across domains, we use four question-answering datasets: (1) ARC2-Easy and (2) ARC2-Challenge (Clark et al., 2018), which are part of the AI2 Reasoning Challenge and require reasoning over grade-school science; and two medical QA benchmarks: (3) MedQA (Jin et al., 2020), consisting of US medical board exam questions, and (4) MedMCQA (Pal et al., 2022), comprising entrance exam questions from the Indian medical school curriculum. All datasets are in multiple-choice format, with ground-truth answers satisfying Assumption A3. Chain-of-Thought reasoning is employed to generate answers. The cost proportion between input and output tokens is set to $\rho=5$, consistent with Anthropic’s current pricing to date (Anthropic, 2025). Details on generation and verification prompts can be found in Appendix B.1. The codebase is available at https://github.com/fanconic/cascaded-llms.
5.1 Cost-Benefit Analysis of Verification Methods
We begin by empirically analysing which verification method from Section 4.1 is most suitable for estimating the confidence of a generated response. Once calibrated via Bayesian logistic regression, these confidence estimates determine whether to defer a prediction from the base model to the larger model.
To assess both cost-efficiency and accuracy, we compare the calibrated base model confidence $\Phi_{\text{base}}$ against two baselines: (1) using only the base model (Llama-1B) and (2) using only the large model (Llama-8B). In Figure 3, we visualise accuracy versus cost per sample across the datasets. We use a threshold-agnostic strategy where deferral to the large model is performed with probability $\Phi_{\text{base}}(x)$. We evaluate four methods: Self-Verification (SV, $n{=}1$), Surrogate Token Probability (STP, $n{=}1$), Consistent Self-Verification (SV, $n{=}5$), and Monte Carlo STP (MC-STP, $n{=}5$). For the latter two, we perform five regenerations or stochastic passes. Each experiment is conducted once using $\mathcal{M}_{\text{base}}$ as the verifying LLM, and once using $\mathcal{M}_{\text{large}}$.
As shown in Figure 3 (and Figures 6, 9, and 12 in Appendix C), using a larger model for verification generally yields a better cost-benefit profile, particularly on simpler datasets (ARC2-Easy, ARC2-Challenge). In contrast, base-model verification provides only marginal gains. On the harder medical datasets (MedQA and MedMCQA), all methods struggle. STP ($n{=}1$) is the most effective in subfigures 3(a) and 3(b).
ARC2-Easy ARC2-Challenge MedQA MedMCQA
Base SV ($n{=}1$) -8.8 $\pm$ 16.4 -13.2 $\pm$ 11.4 -8.0 $\pm$ 9.5 -9.9 $\pm$ 8.9
SV ($n{=}5$) -12.8 $\pm$ 16.5 -19.6 $\pm$ 10.8 -21.3 $\pm$ 8.6 -15.6 $\pm$ 8.7
STP ($n{=}1$) -4.4 $\pm$ 16.7 -5.2 $\pm$ 11.7 -13.6 $\pm$ 9.3 -4.8 $\pm$ 9.3
STP ($n{=}5$) 0.8 $\pm$ 15.2 -9.0 $\pm$ 11.4 -9.4 $\pm$ 9.2 -11.2 $\pm$ 8.8
[…]
SV ($n{=}5$) 50.1 $\pm$ 13.7 10.7 $\pm$ 11.8 -8.7 $\pm$ 9.1 -3.1 $\pm$ 9.0
STP ($n{=}1$) 118.3 $\pm$ 15.8 38.1 $\pm$ 13.7 5.6 $\pm$ 9.7 -3.7 $\pm$ 9.2
STP ($n{=}5$) 97.7 $\pm$ 14.4 15.4 $\pm$ 11.7 -6.8 $\pm$ 9.2 2.6 $\pm$ 9.0
Table 1: $\Delta$IBC Scores for Calibrated Cascade (Llama 1B–8B). Each row indicates a verification method (SV or STP) with $n=1$ or $n=5$, grouped by whether the base or large model was used for verification. A higher score indicates better cost-efficiency relative to the baseline.
To quantitatively assess cost-efficiency, we compute the Incremental Benefit per Cost (IBC) metric from Aggarwal et al. (2024), defined as:
$\text{IBC}_{\text{cascade}}=\frac{P_{\text{cascade}}-P_{\text{base}}}{C_{\phi_{\text{m2m}}}-C_{\text{base}}},\quad\text{IBC}_{\text{base}}=\frac{P_{\text{large}}-P_{\text{base}}}{C_{\text{large}}-C_{\text{base}}},$
where $P$ denotes accuracy and $C$ denotes cost. We then compute the relative gain:
$\Delta\text{IBC}=\frac{\text{IBC}_{{\text{cascade}}}-\text{IBC}_{\text{base}}}{\text{IBC}_{\text{base}}}\cdot 100.$
Higher $\Delta\text{IBC}$ values indicate improved cost-efficiency over the baseline.
As seen in Table 1 verifying with $\mathcal{M}_{\text{large}}$ consistently leads to higher $\Delta$IBC scores, particularly on ARC2-Easy and ARC2-Challenge. On the medical datasets, no single method consistently outperforms the others significantly. Moreover, we see that on the medical datasets, the $\Delta$IBC standard error rates for the verification scores using the large are around 0, indicating no cost-benefit compared to the easier datasets. We report additional results for the other cascades in Tables 2,  3,  4, and 5 in Appendix C. Interestingly, the uncalibrated confidence scores seem to lead to higher $\Delta$IBC, but with a much higher standard error, indicating the instability of uncalibrated confidence scores.
5.2 Online Improvement of the Decision System
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
Figure 4: Cumulative Regret in Online Setting (Llama 1B–8B). Cumulative system risk over time. Training data is collected only when abstentions occur. The cascaded system consistently achieves lower regret. The shaded area indicates the standard error.
Desideratum D3 requires that “the framework should continuously refine its deferral and abstention policies based on feedback, ensuring risk-minimised decision-making”. We simulate an online setting in which the system selects among $\mathcal{M}_{\text{base}}$, $\mathcal{M}_{\text{large}}$, or a human expert, adjusting its thresholds based on feedback from abstentions.
To avoid trivial solutions (e.g., always selecting one model), we balance system risk using $\lambda_{c}=10^{-5}$ and $\lambda_{a}=0.1$, in line with Zellinger et al. (2025). Feedback is collected in batches and stored in a training buffer. Teh deferral and abstention thresholds $\boldsymbol{\theta}=\{\phi_{\text{base}},\xi_{\text{base}},\xi_{\text{large}}\}$ are updated with the ADAM optimiser (Kingma and Ba, 2014), using a learning rate of 0.05 and batch size 10. Thresholds are initialised at $\boldsymbol{\theta}^{(0)}=\{0.5,0.05,0.05\}$, where $\xi_{i}=0.05$ corresponds to the standard deviation of 5% confidence. For the single model baselines, we initialise the abstention threshold als with $\xi=0.05$, and keep the rest of the hyperparameters the same. Throughout this experiment, we use the STP ($n=1$) verification strategy, which was most competitive in the previous section.
We compare the cascaded system $C$ to using only $\mathcal{M}_{\text{base}}$ and only $\mathcal{M}_{\text{large}}$ with a single abstention threshold $\xi$. The system risk for a single model is explained in more detail in Appendix A.1. As we are considering a deployed system, we track the cumulative regret over time, which we define as follows:
$\text{{Regret}}(\mathcal{M})[n]\coloneqq\sum_{t=1}^{n}\mathcal{R}(\mathcal{M}^{(t)}),$
where $\mathcal{M}\in\{C,\mathcal{M}_{\text{base}},\mathcal{M}_{\text{large}}\}$ and $\mathcal{M}^{(t)}$ evolves based on abstention feedback $\mathcal{D}_{t}$.
Figure 4 shows that the cascaded system yields lower cumulative regret over 1000 test samples on ARC2-Easy, ARC2-Challenge, and MedMCQA, compared to using a either model in isolation. On MedQA, gains are less clear, likely due to poor confidence estimation, which were also observed in the section above. Similar trends are observed in other cascades (see Figures  8, 11, 14 in Appendix C). Nevertheless, in three of the four cases, the cascaded LLM system demonstrates lower cumulative regret than when using single models online, where feedback is received when abstained.
6 Limitations
Our cascaded multi-LLM decision-making framework balances accuracy, cost, and abstention but has limitations. Sensitivity to cost and abstention variations can impact efficiency, leading to trivial solutions (only using the cheapest model or the model with the lowest error rate). Discrepancies in model performance or relative costs may cause over-reliance on specific models, reducing adaptability. Furthermore, parameter initialisation affects deferral policy convergence. Additionally, the framework depends on human feedback, which may hinder adaptation if sparse or noisy in a real-world scenario. Finally, fitting a Bayesian logistic regression is usually slightly more complex than a regular one, depending on different posterior approximations or sampling strategies.
7 Conclusion
We proposed a multi-tier decision-making framework that escalates tasks between a base model, a large model, and human experts. By leveraging deferral and abstention policies, our approach aims to enhance performance, accuracy, and abstention while adapting through online learning. Our experiments show that the framework outperforms single-model baselines by reducing unnecessary escalations and improving response correctness on the ARC2-Easy, ARC2-Challenge, and MedMCQA datasets. On MedQA a cascaded model did not outperform the single model approach, potentially due to the complexity of the dataset. Nevertheless, we believe that this proposed system could be beneficial where performance, costs, and abstention of LLMs need to be carefully balanced. Potential applications for this system are in high-stakes environments such as medicine, trading, and education. In this work, we focused on different methods to calculate confidence scores of correct answers, using the posterior predictive’s standard deviation as uncertainty. Future work should investigate different uncertainty quantification methods of LLMs to enhance abstention. Moreover, it would be crucial to examine if there are theoretical guarantees that justify applying cascaded LLMs.
Acknowledgements
We would like to extend our gratitude to Yusuke Kano, Jeremy Voisey, and Alison O’Neil for their insightful discussion and valuable feedback. Claudio Fanconi’s studentship is funded by Canon Medical Systems Corporation.
Appendix A Additional Method Details
A.1 Single-model System Risk
For completeness we report the risk of running either model alone. If $\xi$ is that model’s abstention threshold and $c$ its total cost,
$\mathcal{R}_{\text{single}}(\mathcal{M})=1-\mathbb{E}\!\bigl{[}(1-m_{\text{abst}})\Phi\bigr{]}+\lambda_{c}c+\lambda_{a}\mathbb{E}[m_{\text{abst}}],$
where $m_{\text{abst}}(x)=g_{k}\bigl{(}\Xi(x)-\xi\bigr{)}$.
Appendix B Implementation Details
B.1 Prompts
Throughout this paper, we use prompts to make decision predictions using Chain-of-Thought and verification prompts to determine a response’s factual correctness or uncertainty.
Response Generation Prompt ARC2-Easy + ARC2-Challenge
You are a helpful AI.
Answer the following multiple-choice question using step-by-step reasoning,
then conclude with a final line stating the best answer.
Question: {question}
Choices:
{choice_0}
{choice_1}
{choice_2}
{choice_3}
({choice_4})
Let’s reason step-by-step, then conclude with: "The best answer is: <X>"
Reasoning:
Response Generation Prompt MedQA
You are a medical doctor taking the US Medical Licensing Examination.
Answer the following multiple-choice question using step-by-step reasoning,
then conclude with a final line stating the best answer.
Question: {question}
Choices:
{choice_0}
{choice_1}
{choice_2}
{choice_3}
{choice_4}
Let’s reason step-by-step, then conclude with: "The best answer is: <X>"
Reasoning:
Response Generation Prompt MedMCQA
You are a medical doctor answering real world medical entrance exam questions.
Answer the following multiple-choice question using step-by-step reasoning,
then conclude with a final line stating the best answer.
Question: {question}
Choices:
{choice_0}
{choice_1}
{choice_2}
{choice_3}
Let’s reason step-by-step, then conclude with: "The best answer is: <X>"
Reasoning:
Self Verification Prompt
Given the following question and the model’s answer, please evaluate correctness.
Question: {question}
Model Answer: {candidate_answer}
Please give a confidence score on a scale of 0.0 to 1.0 for this prediction.
Answer:
Surrogate Token Probability Prompt
Given the following question and the model’s answer, please evaluate correctness.
Respond with a single token: {yes_token} or {no_token}
Question: {question}
Model Answer: {candidate_answer}
Is this answer correct: {yes_token} or {no_token}?
Answer:
Appendix C Additional Results
C.1 LLama 1B $\rightarrow$ 8B
ARC2 Easy ARC2 Challenge MedQA MedMCQA
uncal. cal. uncal. cal. uncal. cal. uncal. cal.
Base SV ($n{=}1$) 6.5 $\pm$ 13.1 -8.8 $\pm$ 16.4 2.6 $\pm$ 13.8 -13.2 $\pm$ 11.4 -38.9 $\pm$ 98.5 -8.0 $\pm$ 9.5 -4.2 $\pm$ 12.0 -9.9 $\pm$ 8.9
SV ($n{=}5$) 16.1 $\pm$ 21.3 -12.8 $\pm$ 16.5 -11.3 $\pm$ 18.4 -19.6 $\pm$ 10.8 -55.7 $\pm$ 82.4 -21.3 $\pm$ 8.6 -14.4 $\pm$ 18.6 -15.6 $\pm$ 8.7
STP ($n{=}1$) 6.4 $\pm$ 344.3 -4.4 $\pm$ 16.7 -30.0 $\pm$ 211.7 -5.2 $\pm$ 11.7 -13.9 $\pm$ 87.0 -13.6 $\pm$ 9.3 -21.2 $\pm$ 159.6 -4.8 $\pm$ 9.3
MC-STP ($n{=}5$) -42.5 $\pm$ 116.2 0.8 $\pm$ 15.2 -39.0 $\pm$ 113.6 -9.0 $\pm$ 11.4 -19.8 $\pm$ 107.9 -9.4 $\pm$ 9.2 -42.9 $\pm$ 68.9 -11.2 $\pm$ 8.8
[…]
SV ($n{=}5$) 143.4 $\pm$ 33.6 50.1 $\pm$ 13.7 83.1 $\pm$ 31.6 10.7 $\pm$ 11.8 10.1 $\pm$ 10.2 -8.7 $\pm$ 9.1 63.5 $\pm$ 32.0 -3.1 $\pm$ 9.0
STP ($n{=}1$) 129.4 $\pm$ 19.1 118.3 $\pm$ 15.8 59.1 $\pm$ 16.5 38.1 $\pm$ 13.7 7.3 $\pm$ 9.9 5.6 $\pm$ 9.7 30.0 $\pm$ 11.8 -3.7 $\pm$ 9.2
MC-STP ($n{=}5$) 71.0 $\pm$ 17.2 97.7 $\pm$ 14.4 24.0 $\pm$ 16.5 15.4 $\pm$ 11.7 4.2 $\pm$ 10.0 -6.8 $\pm$ 9.2 9.2 $\pm$ 13.2 2.6 $\pm$ 9.0
Table 2: $\Delta$IBC scores for LLaMA (1B–8B) across datasets and calibration settings. Rows show methods (SV or STP) with $n=1$ or $n=5$, grouped by whether probabilities come from the base or large model. Uncalibrated values are preserved from the original, calibrated values are updated. All values are rounded to 1 decimal place.
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
C.2 Qwen 1.5B $\rightarrow$ 7B
ARC2 Easy ARC2 Challenge MedQA MedMCQA
uncal. cal. uncal. cal. uncal. cal. uncal. cal.
Base SV ($n{=}1$) 278.5 $\pm$ 237.3 -2.4 $\pm$ 32.6 80.5 $\pm$ 174.1 -11.6 $\pm$ 20.5 -45.8 $\pm$ 118.8 -27.3 $\pm$ 14.3 70.4 $\pm$ 89.0 -24.8 $\pm$ 9.5
SV ($n{=}5$) 93.3 $\pm$ 112.7 10.2 $\pm$ 29.6 -15.6 $\pm$ 104.2 -26.2 $\pm$ 19.1 -65.6 $\pm$ 55.3 -25.4 $\pm$ 13.4 29.3 $\pm$ 63.2 -19.5 $\pm$ 9.4
STP ($n{=}1$) -100.0 $\pm$ 948.0 -16.9 $\pm$ 33.2 -100.0 $\pm$ 847.4 -19.3 $\pm$ 20.9 -87.0 $\pm$ 313.1 -23.9 $\pm$ 14.7 -100.0 $\pm$ 835.5 -22.9 $\pm$ 9.7
MC-STP ($n{=}5$) -66.3 $\pm$ 142.6 -37.4 $\pm$ 28.8 -74.3 $\pm$ 143.4 -16.0 $\pm$ 21.0 -69.7 $\pm$ 72.9 -27.3 $\pm$ 13.7 -71.6 $\pm$ 114.7 -22.2 $\pm$ 9.4
[…]
SV ($n{=}5$) 219.7 $\pm$ 35.8 188.5 $\pm$ 31.5 66.5 $\pm$ 22.3 51.0 $\pm$ 23.7 13.9 $\pm$ 18.9 -11.9 $\pm$ 14.7 24.6 $\pm$ 13.8 -11.0 $\pm$ 9.8
STP ($n{=}1$) 284.5 $\pm$ 40.4 242.7 $\pm$ 36.3 96.9 $\pm$ 24.2 89.3 $\pm$ 26.2 21.6 $\pm$ 19.4 1.7 $\pm$ 16.3 35.0 $\pm$ 14.2 -10.0 $\pm$ 10.1
MC-STP ($n{=}5$) 66.5 $\pm$ 19.2 171.4 $\pm$ 36.4 21.3 $\pm$ 14.6 39.2 $\pm$ 22.0 -10.5 $\pm$ 14.6 -16.8 $\pm$ 14.5 -3.5 $\pm$ 10.2 -19.7 $\pm$ 9.3
Table 3: $\Delta$IBC scores for Qwen (1.5B–7B) across datasets and calibration settings. Rows show methods (SV or STP) with $n=1$ or $n=5$, grouped by whether probabilities come from the base or large model. Uncalibrated values are preserved from the original, calibrated values are updated. All values are rounded to 1 decimal place.
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
C.3 LLama 3B $\rightarrow$ 8B
ARC2 Easy ARC2 Challenge MedQA MedMCQA
uncal. cal. uncal. cal. uncal. cal. uncal. cal.
Base SV ($n{=}1$) 267.7 $\pm$ 364.0 3.8 $\pm$ 96.2 322.3 $\pm$ 335.2 -11.5 $\pm$ 81.4 15.7 $\pm$ 205.7 -48.6 $\pm$ 46.9 93.6 $\pm$ 154.5 -27.8 $\pm$ 20.9
SV ($n{=}5$) 247.2 $\pm$ 359.8 32.1 $\pm$ 100.9 352.2 $\pm$ 325.3 0.0 $\pm$ 84.4 12.5 $\pm$ 200.0 -28.2 $\pm$ 49.4 96.1 $\pm$ 146.7 -39.5 $\pm$ 20.4
STP ($n{=}1$) 272.7 $\pm$ 338.6 34.9 $\pm$ 99.6 279.7 $\pm$ 337.5 -40.7 $\pm$ 80.9 -35.6 $\pm$ 199.1 -8.2 $\pm$ 49.3 91.7 $\pm$ 153.0 -19.2 $\pm$ 21.9
MC-STP ($n{=}5$) 272.4 $\pm$ 325.0 44.0 $\pm$ 96.7 279.9 $\pm$ 312.6 33.0 $\pm$ 100.0 -9.2 $\pm$ 204.9 -23.2 $\pm$ 52.8 119.0 $\pm$ 154.3 -45.0 $\pm$ 20.0
[…]
SV ($n{=}5$) 257.4 $\pm$ 324.7 63.2 $\pm$ 106.4 281.6 $\pm$ 292.7 37.4 $\pm$ 83.1 4.7 $\pm$ 131.5 -32.4 $\pm$ 48.8 84.8 $\pm$ 107.0 -14.2 $\pm$ 22.1
STP ($n{=}1$) 331.9 $\pm$ 287.7 77.0 $\pm$ 112.2 254.6 $\pm$ 291.8 -0.7 $\pm$ 83.7 -8.8 $\pm$ 126.8 -34.4 $\pm$ 46.2 95.7 $\pm$ 99.5 -38.6 $\pm$ 20.3
MC-STP ($n{=}5$) 422.2 $\pm$ 328.8 78.2 $\pm$ 109.8 325.5 $\pm$ 306.1 -6.6 $\pm$ 78.7 65.2 $\pm$ 150.7 -52.8 $\pm$ 48.1 91.4 $\pm$ 100.3 -11.5 $\pm$ 22.1
Table 4: $\Delta$IBC scores for LLaMA (3B–8B) across datasets and calibration settings. Rows show methods (SV or STP) with $n=1$ or $n=5$, grouped by whether probabilities come from the base or large model. All values are rounded to 1 decimal place.
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
C.4 Qwen 3B $\rightarrow$ 7B
ARC2 Easy ARC2 Challenge MedQA MedMCQA
uncal. cal. uncal. cal. uncal. cal. uncal. cal.
Base SV ($n{=}1$) 72.6 $\pm$ 67.1 80.6 $\pm$ 98.5 -35.1 $\pm$ 45.8 -3.9 $\pm$ 62.5 -37.5 $\pm$ 27.8 -46.2 $\pm$ 15.7 -21.0 $\pm$ 21.1 -40.3 $\pm$ 14.4
SV ($n{=}5$) 3.7 $\pm$ 48.0 13.7 $\pm$ 76.4 -17.8 $\pm$ 39.0 -59.1 $\pm$ 50.3 -51.3 $\pm$ 20.8 -54.6 $\pm$ 13.2 -39.9 $\pm$ 19.0 -46.2 $\pm$ 13.5
STP ($n{=}1$) -19.4 $\pm$ 23.5 44.7 $\pm$ 97.3 -42.1 $\pm$ 16.8 -53.6 $\pm$ 57.0 -41.5 $\pm$ 12.1 -52.0 $\pm$ 15.5 -37.5 $\pm$ 10.6 -31.5 $\pm$ 15.3
MC-STP ($n{=}5$) -32.1 $\pm$ 21.8 -7.5 $\pm$ 75.5 -43.3 $\pm$ 18.1 -53.2 $\pm$ 53.8 -47.8 $\pm$ 11.9 -50.4 $\pm$ 13.7 -39.9 $\pm$ 11.0 -45.5 $\pm$ 14.0
[…]
SV ($n{=}5$) 99.0 $\pm$ 60.4 139.1 $\pm$ 84.2 21.5 $\pm$ 38.7 27.4 $\pm$ 58.1 -34.9 $\pm$ 14.2 -43.0 $\pm$ 14.2 -13.4 $\pm$ 17.7 -30.3 $\pm$ 14.9
STP ($n{=}1$) 342.9$\pm$ 121.3 313.9 $\pm$ 135.6 96.1 $\pm$ 57.9 99.4 $\pm$ 75.8 -8.3 $\pm$ 20.0 -23.7 $\pm$ 17.4 16.3 $\pm$ 22.2 -17.1 $\pm$ 16.5
MC-STP ($n{=}5$) 21.0 $\pm$ 35.6 97.2 $\pm$ 89.8 -19.4 $\pm$ 24.9 16.2 $\pm$ 56.8 -52.8 $\pm$ 11.1 -44.7 $\pm$ 14.2 -37.4 $\pm$ 12.9 -29.5 $\pm$ 14.9
Table 5: $\Delta$IBC scores for Qwen (3B–7B) across datasets and calibration settings. Rows show methods (SV or STP) with $n=1$ or $n=5$, grouped by whether probabilities come from the base or large model. All values are rounded to 1 decimal place.
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]