### v3

# Cascaded Language Models for Cost-Effective Human–AI Decision-Making

Abstract
A challenge in human-AI decision-making is to balance three factors: the correctness of predictions, the cost of knowledge and reasoning complexity, and the confidence about whether to abstain from automated answers or escalate to human experts. In this work, we present a cascaded LLM decision framework that adaptively delegates tasks across multiple tiers of expertise – a base model for initial candidate answers, a more capable and knowledgeable (but costlier) large model, and a human expert for when the model cascade abstains. Our method proceeds in two stages. First, a deferral policy determines whether to accept the base model’s answer or regenerate it with the large model based on the confidence score. Second, an abstention policy decides whether the cascade model response is sufficiently certain or requires human intervention. Moreover, to overcome static policies and accommodate changing task difficulty, we incorporate an online learning mechanism which uses human feedback. We demonstrate this approach to general question-answering (ARC-Easy, ARC-Challenge, and MMLU) and medical question-answering (MedQA and MedMCQA). Our results demonstrate that our cascaded strategy outperforms single-model baselines in most cases, achieving higher accuracy while reducing costs and providing a principled approach to handling abstentions. [^1] We provide the code for our experiments at https://github.com/fanconic/cascaded-llms
1 Introduction
Data-driven decision support has gained increasing traction in high-stakes fields such as healthcare (Jin et al., 2024; Fan et al., 2024; Li et al., 2024), finance (Li et al., 2023a; Zhao et al., 2024), and education (Xu et al., 2024). For example, in the medical context, large language models (LLMs) can facilitate accurate diagnoses and treatment recommendations that encode vast knowledge Kim et al. (2024). However, high accuracy in such complex settings often requires substantial computational resources or multiple reasoning steps. Additionally, LLMs may hallucinate or generate incorrect outputs with severe consequences. Effective human-AI collaboration should balance correctness, cost, and abstention, ensuring AI-driven assistance integrates seamlessly with expert oversight.
The Challenge.
A key challenge in effective human–AI collaboration is how to allocate computational and human resources efficiently—deciding when an automated model should answer, when it should escalate to a lager and more capable model, and when it should defer to human expertise. Naïve strategies — such as always relying on the cheaper model or always trusting the more capable one — fail to optimise this trade-off. The former increases the risk of errors and hallucinations, while the latter inflates costs. Likewise, static deferral policies, fixed thresholds, or one-off calibrations cannot adapt to changing task distributions or evolving model competence.
These limitations motivate three requirements for any cost-effective human–AI decision-making framework:
  1. 1.
    D1
    Reduce Unnecessary Regenerations: Responses should only be regenerated by a more capable model when there is sufficient evidence that the current one is unreliable..
  2. 2.
    D2
    Abstain when Uncertain: The system should defer to human experts when uncertainty exceeds acceptable bounds, avoiding overconfident automation in high-risk scenarios.
  3. 3.
    D3
    Adapt over time: The framework should continuously refine its deferral and abstention policies as feedback becomes available, ensuring sustained reliability and improvement.
Together, these desiderata define the principles an effective decision-making framework must satisfy, irrespective of implementation.
Our Approach.
We propose a cascaded LLM framework that explicitly satisfies these three requirements. The framework adaptively delegates tasks across multiple tiers of expertise: a lightweight base model provides initial answers; a more capable but costlier large model regenerates responses when confidence is low; and, if uncertainty remains high in the model-generated responses, the system abstains to a human expert. An online learning mechanism continually adjusts the deferral and abstention thresholds based on human feedback, improving decision quality over time. Figure 1 provides an overview of this cascaded decision flow with three example questions of varying difficulty.
Contributions.
Our main contributions are threefold: Cascaded LLM Human-AI Decision System: We introduce a multi-tier decision-making system that coordinates LLMs of varying capacity with human experts to balance accuracy, cost, and abstention. Principled Deferral and Abstention Policies: We design confidence- and uncertainty-based decision policies that regulate when to defer to a larger model or abstain to humans, guided by Bayesian calibration for reliable verification. Online Learning for Adaptive Decision-Making: We propose an online optimisation scheme that refines the deferral and abstention thresholds using human feedback, enabling continual adaptation to task complexity.
Figure 1: Cascaded LLM Human-AI Decision-Making Framework Examples. Given a decision-making problem, the system (1) generates an initial response with a base model, (2) verifies correctness probability, (2.5) defers to a larger model if needed, (3) assesses response uncertainty, and (3.5) abstains to a human expert if necessary. If feedback is available, deferral and abstention modules are adjusted over time. For this system to work efficiently, the modules should uphold three desiderata: D1 the deferral policy regenerates responses only when necessary, D2 the abstention policy escalates to humans only when uncertainty is high, D3 the system continuously improves with feedback.
2 Related Work
Multi-LLM Answer Generation. Several studies have explored collaborative frameworks that leverage multiple LLMs of varying capacities to enhance both performance and cost-efficiency beyond the capabilities of a single model (Chen et al., 2023; Ding et al., 2024; Aggarwal et al., 2024). Chen et al. (2023) proposed cost-effective strategies such as prompt structuring, model approximation, and cascaded LLM frameworks. Similarly, Ding et al. (2024) introduced an intelligent routing mechanism that dynamically assigns prompts to the most appropriate model. Aggarwal et al. (2024) developed a black-box LLM framework for cost-efficient response generation, formalised as a Partially Observable Markov Decision Process (POMDP), requiring minimal training data. Zhu et al. (2023a) proposed a multiplexer-based approach that balances queries between a small and a large LLM, employing a trained BERT classifier to determine when the smaller model suffices. Šakota et al. (2024) introduced a meta-model-driven selection framework that requires pre-training for optimal query distribution. In a parallel line of research, speculative decoding (Leviathan et al., 2023), employs a lightweight model to generate multiple tokens, which a larger model subsequently verifies.
In contrast to prior research, we propose a multi-tier framework for human-AI collaboration. Rather than relying solely on automation, our approach integrates human intervention when model uncertainty is too high, addressing a gap in previous multi-tier frameworks. Compared to speculative decoding research, our work prioritises the factual correctness of complete responses rather than token-wise distributions, enabling more robust decision-making rather than just fluent text generation. Zellinger et al. (2025) conducts a concurrent line of research that is closest to our work on cascaded LLMs, as well as in their previous works (Zellinger and Thomson, 2024, 2025). They focus on probabilistic modelling of cascading LLMs and their deferral and abstention mechanisms.
LLM Answer Verification and Uncertainty Quantification. Ensuring the reliability of LLM-generated responses requires adequate verification and uncertainty quantification mechanisms. Several studies have explored self-verification strategies (Weng et al., 2023; Jiang et al., 2024; Pan et al., 2024), often leveraging the LLM’s internal knowledge (Dhuliawala et al., 2023). Alternative approaches employ external knowledge sources for verification (Pan et al., 2024; Gao et al., 2023; Peng et al., 2023). Aggarwal et al. (2024) introduced verification techniques based on available contextual information, predominantly involving multiple LLM queries to validate response accuracy. Another research direction quantifies factual correctness uncertainty (Mahaut et al., 2024). Kadavath et al. (2022) conducted a detailed analysis of how LLMs express uncertainty through surrogate token probabilities, demonstrating their effectiveness in calibration. Azaria and Mitchell (2023) explored internal LLM states, training classifiers to quantify uncertainty, while methods such as semantic uncertainty estimation (Kuhn et al., 2023) enhance robustness by analysing variations in semantically equivalent token sequences.
Our approach relies on surrogate token probability (Kadavath et al., 2022) as a core verification component. However, we extend this methodology by integrating a hierarchical escalation mechanism that dynamically transitions between models and human experts based on verification results.
Selective Prediction. Selective prediction enables models to abstain from uncertain queries (El-Yaniv and Wiener, 2010), a crucial feature in risk-sensitive settings where errors are costly. The idea dates back to Chow’s work on optical character recognition (Chow, 1957, 1970), and has since been shown to improve deep learning performance (Geifman and El-Yaniv, 2017). In NLP, abstention has been introduced through confidence-based thresholds (Xin et al., 2021; Yoshikawa and Okazaki, 2023), with recent work on uncertainty quantification for large language models advancing this line of research (Manakul et al., 2023; Farquhar et al., 2024; Lin et al., 2024).
LLMs in Online Learning. Traditional LLM research predominantly evaluates language models on static datasets. However, our work aligns with online learning paradigms, wherein policies are continuously refined in response to streaming data (Cortes et al., 2018; Ye et al., 2024). Our methodology is inspired by Jarrett et al. (2022), who introduced an online decision mediation framework mediating between suboptimal human decisions and an expert oracle. A similar research with the online learning approach is conducted by Zhu et al. (2023a), which extended their multiplexer mechanism to an online setting.
3 Background
3.1 Cascaded Decision System
We consider a two-tiered cascaded LLM decision system for question answering under resource constraints, denoted by $C=\mathcal{M}_{\text{base}}\rightarrow\mathcal{M}_{\text{large}}$, following the notation of Zellinger et al. (2025). Let $x\in\mathcal{X}$ be a problem statement or prompt, and let $y\in\mathcal{Y}$ denote a system-generated response. For every input $x$, the models return a confidence score $\Phi_{i}(x)\in[0,1]$ and an uncertainty score $\Xi_{i}(x)\in[0,\infty)$, where $i\in\{\text{base},\text{large}\}$. The decision to predict using the base model $\mathcal{M}_{\text{base}}$ or to defer to the larger model $\mathcal{M}_{\text{large}}$ is based on whether the confidence exceeds a deferral threshold, i.e., $\Phi(x)>\phi_{\text{base}}$. Thus, a prediction is only made if the base model is sufficiently confident. In contrast, abstention is governed by predictive uncertainty: if this exceeds a threshold, $\Xi_{i}(x)>\xi_{i}$, the system abstains and forwards the query to a human expert.
We formally define the cascaded decision system as:
[…]
The decision flow of this cascade is also illustrated in Figure 2. While we focus on a two-model system here, the framework naturally generalises to cascades involving multiple LLMs of varying sizes.
The objective is to generate accurate responses while accounting for the computational costs of the models and abstaining when the system is too uncertain. As described in Zellinger and Thomson (2024), this constitutes a multi-objective optimisation problem over three dimensions: error, cost, and abstention. Formally, we minimise the system risk:
$\mathcal{R}(C)=\mathbb{P}(\text{error}\land\neg\text{abstention})+\lambda_{c}\mathbb{E}[\text{Cost}]+\lambda_{a}\mathbb{P}(\text{abstention})$ (2)
Here, $\mathbb{P}(\text{error}\land\neg\text{abstention})$ denotes the probability of the system making an error when it does not abstain, $\mathbb{E}[\text{Cost}]$ is the expected computational cost incurred, and $\mathbb{P}(\text{abstention})$ is the probability of the system abstaining and deferring to a human expert. The terms $\lambda_{c}$ and $\lambda_{a}$ weight the cost and abstention penalties, respectively. We explain the system risk in more detail in Section 4.2.
Assumptions.
A1
The base model is cost-efficient but less accurate, whereas the large model is more capable but computationally expensive. A2 Generating responses incurs significantly higher cost than processing inputs, especially in settings that require Chain-of-Thought (CoT) prompting (Wei et al., 2022) or advanced test-time reasoning (Xie et al., 2024). A3 Each response is assumed to be either correct or incorrect, with no ambiguity.
Figure 2: Decision flow of the two-tiered cascaded LLM system. The base model first evaluates each query. Confident, low-uncertainty responses are accepted; uncertain ones are passed to the large model or, if still uncertain, deferred to a human expert. Online feedback progressively improves these policies.
3.2 Cost Calculation
To estimate the computational cost of response generation, we define a cost function that scales linearly with model size and token counts. Let $s$ denote the model size (in billions of parameters, e.g., Llama3.1-8B $\Rightarrow s=8$), $t_{\text{in}}$ and $t_{\text{out}}$ the numbers of input and output tokens, and $\rho>0$ the output-to-input token cost ratio accounting for the higher cost of generation. The total cost is given by:
$\text{Cost}(s,t_{\text{in}},t_{\text{out}},\rho)=s\cdot\left(t_{\text{in}}+\rho\cdot t_{\text{out}}\right)$ (3)
This provides a simple yet effective way to compare models of different sizes under a unified cost metric, independent of infrastructure-specific pricing. Additional cost components can be incorporated as needed.
4 Methods
4.1 Calibrated Confidence and Uncertainty Estimation
Effective deferral and abstention decisions in a cascaded system critically depend on accurately quantifying model confidence $\Phi(x)$ and uncertainty $\Xi(x)$ for each input $x$. Overconfident or miscalibrated predictions can lead to errors, while excessive uncertainty may result in unnecessary escalations. Therefore, the first part of our method focuses on analysing a range of techniques to estimate these quantities in a reliable and cost-efficient manner. To this end, we evaluate four complementary methods that approximate the probability that a response is correct.
(1) Self-Verification.
Given an input $x$ and a model response $y_{\text{base}}\sim\mathcal{M}_{i}(x)$, we prompt the same model to quantify how likely the response is correct by generating a new response (Li et al., 2023b). The model returns raw confidence score by outputting either a scalar value token in response to a verification prompt (see Appendix B.2). The outputted probability serves as an uncalibrated estimate of correctness.
(2) Consistent Self-Verification.
We repeat the self-verification process $n$ times under stochastic sampling (e.g., with temperature), and aggregate the resulting probabilities. The empirical mean forms the uncalibrated confidence score. This approach is inspired by self-consistency as in (Aggarwal et al., 2024).
(3) Surrogate Token Probability.
We adopt the approach of Kadavath et al. (2022), where the model $\mathcal{M}_{i}$ is asked to verify whether a generated response $y$ is correct, and we extract the next-token probability over the discrete label set YES/NO. Specifically:
$p_{i}(x)=\frac{\mathcal{M}_{i}(\texttt{YES}\mid x,y)}{\mathcal{M}_{i}(\texttt{YES}\mid x,y)+\mathcal{M}_{i}(\texttt{NO}\mid x,y)},$ (4)
(4) Monte-Carlo Surrogate Token Probability.
To obtain better confidence estimates, we apply Monte Carlo Dropout (Gal and Ghahramani, ) at test time when computing the surrogate token probability. For each of $n$ stochastic forward passes, we sample an estimate $\hat{p}^{(t)}_{i}(x)$, and the empirical mean forms the uncalibrated confidence score:
${p}_{i}(x)=\frac{1}{T}\sum_{t=1}^{n}\hat{p}^{(t)}_{i}(x)$ (5)
Model Evaluation by Larger Models.
For each of the above methods, the evaluating model $\mathcal{M}_{i}$ can either be the same model that generated the original response, or a larger model in the cascade, if available. While self-evaluation is cheap and self-contained, verifying a small model’s output using a larger model is still substantially cheaper than generating a new response from scratch—particularly when generation involves long-form reasoning, as per Assumption A2. Additionally, larger models tend to be better calibrated and may yield more reliable verification, improving downstream deferral and abstention decisions (Zhu et al., 2023b; Chhikara, 2025).
Bayesian Calibration.
To ensure that the extracted confidence scores are comparable across models and consistent with empirical correctness, we fit a Bayesian logistic regression model on a small calibration set of 100 samples. This is a Bayesian version of Platt scaling (Platt, 2000), and we assume a Normal distribution as prior. We follow Zellinger and Thomson (2024)’s approach and apply a non-linear transformation on the raw confidence score before inputting it into the Bayesian model, to spread out the clusters of overconfident probabilities.
$p_{tr}(p_{i})=\begin{cases}\log(\frac{1}{1-p_{i}})&\text{if }p_{i}\geq 0.5\\ \log(2)-\log(\frac{1}{p_{i}})&\text{if }p_{i}<0.5\end{cases}$ (6)
Subsequently, the Bayesian Logistic Regression outputs a posterior distribution over correctness. The mean of the posterior predictive distribution defines the calibrated confidence $\Phi(x)$, while we use standard deviation as a model-based uncertainty estimate $\Xi(x)$, as in (Fanconi et al., 2023).
4.2 Online Improvement
To enable online learning D3, we parameterise the deferral and abstention thresholds and optimise them online. Given a dataset $\mathcal{D}^{(t)}$ at time $t\in\mathbb{N}$ with previous problem statements and ground truth labels, we update the thresholds using stochastic gradient descent. While the system is deployed, we assume that we will receive a ground truth response ($y^{*}$) at the end of every decision if the system abstains. Thus, our dataset continually increases $\mathcal{D}^{(t)}=\mathcal{D}^{(t-1)}\cup\{x,y^{*}\}$ every time the cascade abstains.
Our objective function is the system risk $\mathcal{R}(C)$ (Equation 2). We expand this risk into the concrete, differentiable losses. Throughout, let
$\Phi_{i}(x)\in[0,1],\qquad\Xi_{i}(x)\in[0,1],\qquad i\in\{\text{base},\text{large}\}$
denote the calibrated probability of correctness (posterior predictive) and a uncertainty score (i.e. posterior predictive standard deviation) returned by model $i$ for an input $x$. The optimisation variables are
$\phi_{\text{base}},\;\xi_{\text{base}},\;\xi_{\text{large}}\in(0,1),$
For numerical stability we treat their raw, unconstrained versions $\phi_{\text{base}}^{\text{raw}},\xi_{\text{base}}^{\text{raw}},\xi_{\text{large}}^{\text{raw}}\in\mathbb{R}$ as the true optimisation parameters and map them to $(0,1)$ with a sigmoid function:
$\phi_{\text{base}}=\sigma\!\bigl(\phi_{\text{base}}^{\text{raw}}\bigr),\quad\xi_{\text{base}}=\sigma\!\bigl(\tau_{\text{base}}^{\text{raw}}\bigr),\quad\xi_{\text{large}}=\sigma\!\bigl(\tau_{\text{large}}^{\text{raw}}\bigr).$
To keep the loss fully differentiable, we replace every Boolean test with a soft logistic step, where $k$ determines the steepness
$\mathbf{1}\{z>0\}\;\longrightarrow\;g_{k}(z)=\sigma(k\,z).$
With this convention the three mutually exclusive masks at the base stage are
$p_{\text{abst1}}(x)$ $=g_{k}\!\bigl(\Xi_{\text{base}}(x)-\xi_{\text{base}}\bigr),$ (7)
$m_{\text{pred1}}(x)$ $=\bigl(1-m_{\text{abst1}}\bigr)\cdot g_{k}\!\bigl(\Phi_{\text{base}}(x)-\phi_{\text{base}}\bigr),$ (8)
$m_{\text{defer1}}(x)$ $=\bigl(1-m_{\text{abst1}}\bigr)\cdot g_{k}\!\bigl(\phi_{\text{base}}-\Phi_{\text{base}}(x)\bigr),$ (9)
and the masks at the large stage are
$p_{\text{abst2}}(x)$ $=m_{\text{defer1}}(x)\cdot g_{k}\!\bigl(\Xi_{\text{large}}(x)-\xi_{\text{large}}\bigr),$ (10)
$m_{\text{pred2}}(x)$ $=m_{\text{defer1}}(x)\cdot\bigl(1-g_{k}(\Xi_{\text{large}}(x)-\xi_{\text{large}})\bigr).$ (11)
Probability of abstention. The cascade abstains in two mutually exclusive ways, so
$\mathbb{P}(\text{abstention})\;=\;p_{\text{abst1}}\;+{p_{\text{abst2}}}.$ (12)
Expected correctness. Only the prediction masks contribute a non-zero probability of correctness; we weight each by the calibrated confidence:
$\mathbb{E}[\text{Correct}]\;=\;\mathbb{E}\!\bigl[m_{\text{pred1}}\cdot\Phi_{\text{base}}\bigr]\;+\;\mathbb{E}\!\bigl[m_{\text{pred2}}\cdot\Phi_{\text{large}}\bigr].$ (13)
Expected cost. Let $c_{1}$ be the costs from the base model, which consist of the generation cost and the verification cost (either by itself or by a larger model). Furthermore, $c_{2}$ is the generation cost and the verification cost caused by the large model. The first term is incurred on every query; the second is incurred only if we defer:
$\mathbb{E}[\text{Cost}]\;=\;c_{1}\;+\;\mathbb{E}[m_{\text{defer1}}]\cdot c_{2}.$ (14)
System-risk objective. Substituting the three expectations above into Eq. (2) produces the differentiable loss that is back-propagated during threshold optimisation in online learning:
$\mathcal{R}(C)$ $=1-\mathbb{E}[\text{Correct}]\;+\;\lambda_{c}\,\mathbb{E}[\text{Cost}]\;+\;\lambda_{a}\,\bigl(p_{\text{abst1}}+p_{\text{abst2}}\bigr).$ (15)
5 Experiments
In this section, we empirically assess whether the desiderata D1, D2, and D3, introduced in Section 1, are satisfied. For D1 and D2, we analyse in Section 5.1 the performance of various confidence estimation techniques with respect to calibration and cost-efficiency. Subsequently, in Section 5.2, we investigate whether the system improves through online learning.
General Setup.
[…]
To evaluate the generalisability of our framework across domains, we use five question-answering datasets: (1) ARC2-Easy and (2) ARC2-Challenge (Clark et al., 2018), which are part of the AI2 Reasoning Challenge and require reasoning over grade-school science; (3) Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2021), which covers 57 subjects ranging from complex STEM to international law, nutrition, and religion; and two medical QA benchmarks: (4) MedQA (Jin et al., 2020), consisting of US medical board exam questions, and (5) MedMCQA (Pal et al., 2022), comprising entrance exam questions from the Indian medical school curriculum. All datasets are in multiple-choice format, with ground-truth answers satisfying Assumption A3. Chain-of-Thought reasoning is employed to generate answers. The cost proportion between input and output tokens is set to $\rho=5$, consistent with Anthropic’s current pricing to date (Anthropic, 2025). Details on generation and verification prompts can be found in Appendix B.2.
5.1 Cost-Benefit Analysis of Verification Methods
We begin by empirically analysing which verification method from Section 4.1 is most suitable for estimating the confidence of a generated response. Once calibrated via Bayesian logistic regression, these confidence estimates determine whether to defer a prediction from the base model to the larger model.
To assess both cost-efficiency and accuracy, we compare the calibrated base model confidence $\Phi_{\text{base}}$ against two baselines: (1) using only the base model (Qwen-2.5-1.5B) and (2) using only the large model (Qwen-2.5-7B). In Figure 3, we visualise accuracy versus cost per sample across the datasets. We use a threshold-agnostic strategy where deferral to the large model is performed with probability $\Phi_{\text{base}}(x)$. We evaluate four methods: Self-Verification (SV, n=1n{=}1), Surrogate Token Probability (STP, n=1n{=}1), Consistent Self-Verification (SV, n=5n{=}5), and Monte Carlo STP (MC-STP, n=5n{=}5). For the latter two, we perform five regenerations or stochastic passes. Each experiment is conducted once using $\mathcal{M}_{\text{base}}$ as the verifying LLM, and once using $\mathcal{M}_{\text{large}}$.
[…]
As shown in Figure 3 (and Figures 7, 10, and 13 in Appendix C), using a larger model for verification generally yields a better cost-benefit profile, particularly on simpler datasets (ARC2-Easy, ARC2-Challenge, MMLU). In contrast, base-model verification provides only marginal gains. On the more complex medical datasets (MedQA and MedMCQA), all methods struggle. STP (n=1n{=}1) is the most effective the ARC2-Easy and ARC2-Challenge dataset.
To quantitatively assess cost-efficiency, we compute the Incremental Benefit per Cost (IBC) metric from Aggarwal et al. (2024), defined as:
$\text{IBC}_{\text{cascade}}=\frac{P_{\text{cascade}}-P_{\text{base}}}{C_{\phi_{\text{m2m}}}-C_{\text{base}}},\quad\text{IBC}_{\text{base}}=\frac{P_{\text{large}}-P_{\text{base}}}{C_{\text{large}}-C_{\text{base}}},$
where $P$ denotes accuracy and $C$ denotes cost. We then compute the relative gain:
$\Delta\text{IBC}=\frac{\text{IBC}_{{\text{cascade}}}-\text{IBC}_{\text{base}}}{\text{IBC}_{\text{base}}}\cdot 100.$
Higher $\Delta\text{IBC}$ values indicate improved cost-efficiency over the baseline.
ARC2 Easy ARC2 Challenge MMLU MedQA MedMCQA
Base SV (n=1n{=}1) -2.4 $\pm$ 32.6 -11.6 $\pm$ 20.5 -18.7 $\pm$ 4.9 -27.3 $\pm$ 14.3 -24.8 $\pm$ 9.5
SV (n=5n{=}5) 10.2 $\pm$ 29.6 -26.2 $\pm$ 19.1 -19.4 $\pm$ 4.9 -25.4 $\pm$ 13.4 -19.5 $\pm$ 9.4
STP (n=1n{=}1) -16.9 $\pm$ 33.2 -19.3 $\pm$ 20.9 -20.4 $\pm$ 4.5 -23.9 $\pm$ 14.7 -22.9 $\pm$ 9.7
MC-STP (n=5n{=}5) -37.4 $\pm$ 28.8 -16.0 $\pm$ 21.0 -24.6 $\pm$ 4.2 -27.3 $\pm$ 13.7 -22.2 $\pm$ 9.4
Large SV (n=1n{=}1) 258.4 $\pm$ 39.1 70.5 $\pm$ 26.0 -7.8 $\pm$ 4.7 -6.1 $\pm$ 15.8 -11.6 $\pm$ 10.0
SV (n=5n{=}5) 188.5 $\pm$ 31.5 51.0 $\pm$ 23.7 -24.1 $\pm$ 3.9 -11.9 $\pm$ 14.7 -11.0 $\pm$ 9.8
STP (n=1n{=}1) 242.7 $\pm$ 36.3 89.3 $\pm$ 26.2 2.5 $\pm$ 5.2 1.7 $\pm$ 16.3 -10.0 $\pm$ 10.1
MC-STP (n=5n{=}5) 171.4 $\pm$ 36.4 39.2 $\pm$ 22.0 -22.4 $\pm$ 4.0 -16.8 $\pm$ 14.5 -19.7 $\pm$ 9.3
[…]
5.2 Online Improvement of the Decision System
Desideratum D3 requires that “the framework should continuously refine its deferral and abstention policies as feedback becomes available, ensuring sustained reliability and improvement”. We simulate an online setting in which the system selects among $\mathcal{M}_{\text{base}}$, $\mathcal{M}_{\text{large}}$, or a human expert, adjusting its thresholds based on feedback from abstentions.
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MMLU
(d) MedQA
(e) MedMCQA
Figure 4: Cumulative Regret in Online Setting (Qwen-2.5 1.5B $\rightarrow$ 7B). Cumulative system risk over time. Training data is collected only when abstentions occur. The cascaded system consistently achieves lower regret.
Specifically, the experiment streams 1,000 unseen questions in a random order to simulate production traffic. Before the first query, the deferral probabilities are calibrated using the values learned from the 100-sample calibration set. Thereafter, we add queries that were marked for abstention and answered by an oracle expert to a replay buffer. This replay buffer is used to perform the ADAM optimiser (Kingma and Ba, 2014) updates on the differentiable risk 2 with a learning rate of 0.05 and a batch size of 10, on the deferral and abstention thresholds $\boldsymbol{\theta}=\{\phi_{\text{base}},\xi_{\text{base}},\xi_{\text{large}}\}$. The prediction is made on an unseen query and the regret is calculated on it. If a query is added to the replay buffer, the regret associated with it has already been calculated, and the sample now becomes a training sample; however, it is not further evaluated. We compare the cascaded system $C$ to using only $\mathcal{M}_{\text{base}}$ and only $\mathcal{M}_{\text{large}}$ with a single abstention threshold $\xi$. The system risk for a single model is explained in more detail in Appendix A.1. As we are considering a deployed system, we track the cumulative regret over time, which we define as follows:
$\text{{Regret}}(\mathcal{M})[n]\coloneqq\sum_{t=1}^{n}\mathcal{R}(\mathcal{M}^{(t)}),$
where $\mathcal{M}\in\{C,\mathcal{M}_{\text{base}},\mathcal{M}_{\text{large}}\}$ and $\mathcal{M}^{(t)}$ evolves based on abstention feedback $\mathcal{D}_{t}$. We chose regret as the metric for this experiment, inspired by the work on online decision mediation (Jarrett et al., 2022). Regret is the running sum of our per‑query risk. Because error, compute, and human‑hand‑off are already weighted into the same units, adding them over time tells you the exact “bill” the system has paid. A lower regret curve indicates a higher benefit, as it represents a combination of abstentions, correct predictions, and costs, and we can see it grow over time in a deployed setting. The regret curve illustrates how quickly a policy learns online and whether early mistakes are compensated for later.
We initialise the thresholds at $\boldsymbol{\theta}^{(0)}=\{0.5,0.05,0.05\}$, where $\xi_{i}=0.05$ corresponds to the standard deviation of 5% confidence. For the single model baselines, we initialise the abstention threshold as with $\xi=0.05$, and keep the rest of the hyperparameters the same. Throughout this experiment, we employ the STP ($n=1$) verification strategy, which was found to be the most competitive in the previous section. To avoid trivial solutions (e.g., always selecting one model), we balance system risk using $\lambda_{c}=10^{-5}$ and $\lambda_{a}=0.1$, in line with Zellinger et al. (2025).
Figure 4 shows that the cascaded system yields lower cumulative regret over 1000 test samples on ARC2-Easy, ARC2-Challenge, MMLU, and MedMCQA, compared to using either model in isolation. On MedQA, gains are less clear, likely due to poor confidence estimation, which was also observed in the section above. Similar trends are observed in other cascades (see Figures 9, 12, 15 in Appendix C). Nevertheless, in four of the five cases, the cascaded LLM system demonstrates lower cumulative regret than when using single models online, where feedback is received when abstaining from action. Additionally, we experiment by comparing our proposed gradient-based approach to a traditional grid search over $\mathbf{\theta}$ in Appendix C.7.
5.3 The Effect of Imperfect Expert
Figure 5: Imperfect Experts. We increase the percentage of flipped labels during system calibration, simulating imperfect experts, which in turn increases the system’s risk of error.
To recall the system risk objective in Equation 15, the part where wrong human annotations will have an impact is the expected correctness (Equation 13). More precisely, the calibrated confidence scores $\Phi_{\text{base}}$ and $\Phi_{\text{large}}$. The noisier the feedback is, the more uncalibrated $\Phi_{\text{base}}$ and $\Phi_{\text{large}}$ will become. Therefore, optimisation of the cascaded model will become unreliable.
We demonstrate this through an additional experiments in the online setting, where we progressively swap correct to incorrect predictions while calibrating the model, and how this affects the trajectory on the ARC-Easy dataset with Qwen-2.5-1.5B $\rightarrow$ Qwen-2.5-7B. The results are displayed in Figure 5. We observe that the higher the percentage of label corruption is in the calibration set, the higher the cumulative regret becomes while deploying the decision-making system.
6 Limitations
Our cascaded multi-LLM decision-making framework strikes a balance between accuracy, cost, and abstention, but it has limitations. Sensitivity to cost and abstention variations can impact efficiency, leading to trivial solutions (only using the cheapest model or the model with the lowest error rate). Discrepancies in model performance or relative costs may lead to over-reliance on specific models, thereby reducing adaptability. Furthermore, parameter initialisation affects the convergence of the deferral policy. Additionally, the framework relies on human feedback, which may hinder adaptation if it is sparse or noisy in a real-world scenario. Finally, fitting a Bayesian logistic regression model is usually more complex than fitting a regular one, depending on the different posterior approximations or sampling strategies employed.
7 Conclusion
We proposed a multi-tier decision-making framework that escalates tasks between a base model, a large model, and human experts. By leveraging deferral and abstention policies, our approach aims to enhance performance, accuracy, and abstention while adapting through online learning. Our experiments show that the framework outperforms single-model baselines by reducing unnecessary escalations and improving response correctness on the ARC2-Easy, ARC2-Challenge, MMLU, and MedMCQA datasets. On MedQA, a cascaded model did not outperform the single model approach, potentially due to the complexity of the dataset. Nevertheless, we believe that this proposed system could be beneficial where performance, costs, and abstention of LLMs need to be carefully balanced. Future work should investigate different uncertainty quantification methods of LLMs to enhance abstention. Moreover, it would be crucial to examine whether there are theoretical guarantees that justify the application of cascaded LLMs.
Acknowledgements and Disclosure of Funding
We want to extend our gratitude to Alan Jeffares, Paulius Rauba, Yusuke Kano, Jeremy Voisey, and Alison Smithard for their insightful discussions and valuable feedback. Canon Medical Systems Corporation funds CF’s studentship. This work was supported by Microsoft’s Accelerate Foundation Models Academic Research initiative.
Appendix A Additional Method Details
A.1 Single-model System Risk
For completeness we report the risk of running either model alone. If $\xi$ is that model’s abstention threshold and $c$ its total cost,
$\mathcal{R}_{\text{single}}(\mathcal{M})=1-\mathbb{E}\!\bigl[(1-m_{\text{abst}})\Phi\bigr]+\lambda_{c}c+\lambda_{a}\mathbb{E}[m_{\text{abst}}],$
where $m_{\text{abst}}(x)=g_{k}\bigl(\Xi(x)-\xi\bigr)$.
Appendix B Implementation Details
The code for this paper, to reproduce the results is provided at https://github.com/fanconic/cascaded-llms. All experiments are implemented in Python [Van Rossum and Drake Jr, 1995] with PyTorch [Paszke et al., 2017] and Hugging Face Transformers [Wolf et al., 2020].
Compute.
Experiments are conducted on a single A100-class GPUs.
B.1 Generation Models
Policies are initialised from instruction-tuned checkpoints and trained with the learned reward signal. The following policy backbones are used:
B.2 Prompts
Throughout this paper, we use prompts to make decision predictions using Chain-of-Thought and verification prompts to determine a response’s factual correctness or uncertainty.
Response Generation Prompt ARC2-Easy + ARC2-Challenge
You are a helpful AI.
Answer the following multiple-choice question using step-by-step reasoning,
then conclude with a final line stating the best answer.
Question: {question}
Choices:
{choice_0}
{choice_1}
{choice_2}
{choice_3}
({choice_4})
Let’s reason step-by-step, then conclude with: "The best answer is: <X>"
Reasoning:
Response Generation Prompt MMLU
You are an expert in {subject}.
Answer the following multiple-choice question using step-by-step reasoning,
then conclude with a final line stating the best answer.
Question: {question}
Choices:
{choices}
Let’s reason step-by-step, then conclude with: "The best answer is: <X>"
Reasoning:
Response Generation Prompt MedQA
You are a medical doctor taking the US Medical Licensing Examination.
Answer the following multiple-choice question using step-by-step reasoning,
then conclude with a final line stating the best answer.
Question: {question}
Choices:
{choice_0}
{choice_1}
{choice_2}
{choice_3}
{choice_4}
Let’s reason step-by-step, then conclude with: "The best answer is: <X>"
Reasoning:
Response Generation Prompt MedMCQA
You are a medical doctor answering real world medical entrance exam questions.
Answer the following multiple-choice question using step-by-step reasoning,
then conclude with a final line stating the best answer.
Question: {question}
Choices:
{choice_0}
{choice_1}
{choice_2}
{choice_3}
Let’s reason step-by-step, then conclude with: "The best answer is: <X>"
Reasoning:
Self Verification Prompt
Given the following question and the model’s answer, please evaluate correctness.
Question: {question}
Model Answer: {candidate_answer}
Please give a confidence score on a scale of 0.0 to 1.0 for this prediction.
Answer:
Surrogate Token Probability Prompt
Given the following question and the model’s answer, please evaluate correctness.
Respond with a single token: {yes_token} or {no_token}
Question: {question}
Model Answer: {candidate_answer}
Is this answer correct: {yes_token} or {no_token}?
Answer:
Appendix C Additional Results
C.1 Qwen-2.5 1.5B $\rightarrow$ 7B
C.1.1 Uncalibrated
Figure 6: Benefit-Cost Analysis of Uncalibrated Verification Methods (Qwen-2.5 1.5B-7B). We display the cost vs accuracy of the various verification methods, using the cascade ($\texttt{Qwen-2.5-1.5B}\rightarrow\texttt{Qwen-2.5-7B})$. Verification methods, which are located above the linear interpolation between the base or large models, indicate a positive cost-benefit ratio. The error bars indicate the standard error.
ARC2 Easy ARC2 Challenge MMLU MedQA MedMCQA
Base SV (n=1n{=}1) 278.5 $\pm$ 237.3 80.5 $\pm$ 174.1 -18.7 $\pm$ 4.9 -45.8 $\pm$ 118.8 70.4 $\pm$ 89.0
SV (n=5n{=}5) 93.3 $\pm$ 112.7 -15.6 $\pm$ 104.2 -19.4 $\pm$ 4.9 -65.6 $\pm$ 55.3 29.3 $\pm$ 63.2
STP (n=1n{=}1) -100.0 $\pm$ 948.0 -100.0 $\pm$ 847.4 -20.4 $\pm$ 4.5 -87.0 $\pm$ 313.1 -100.0 $\pm$ 835.5
MC-STP (n=5n{=}5) -66.3 $\pm$ 142.6 -74.3 $\pm$ 143.4 -24.6 $\pm$ 4.2 -69.7 $\pm$ 72.9 -71.6 $\pm$ 114.7
Large SV (n=1n{=}1) 280.5 $\pm$ 41.6 84.8 $\pm$ 24.6 -3.0 $\pm$ 5.0 24.7 $\pm$ 21.0 32.0 $\pm$ 14.0
SV (n=5n{=}5) 219.7 $\pm$ 35.8 66.5 $\pm$ 22.3 -7.0 $\pm$ 4.8 13.9 $\pm$ 18.9 24.6 $\pm$ 13.8
[…]
MC-STP (n=5n{=}5) 66.5 $\pm$ 19.2 21.3 $\pm$ 14.6 -5.7 $\pm$ 4.9 -10.5 $\pm$ 14.6 -3.5 $\pm$ 10.2
[…]
C.2 LLama3 1B $\rightarrow$ 8B
C.2.1 Calibrated
[…]
ARC2 Easy ARC2 Challenge MMLU MedQA MedMCQA
Base SV (n=1n{=}1) -8.8 $\pm$ 16.4 -13.2 $\pm$ 11.4 -9.6 $\pm$ 2.6 -8.0 $\pm$ 9.5 -9.9 $\pm$ 8.9
SV (n=5n{=}5) -12.8 $\pm$ 16.5 -19.6 $\pm$ 10.8 -11.9 $\pm$ 2.6 -21.3 $\pm$ 8.6 -15.6 $\pm$ 8.7
STP (n=1n{=}1) -4.4 $\pm$ 16.7 -5.2 $\pm$ 11.7 -6.8 $\pm$ 2.8 -13.6 $\pm$ 9.3 -4.8 $\pm$ 9.3
MC-STP (n=5n{=}5) 0.8 $\pm$ 15.2 -9.0 $\pm$ 11.4 -9.0 $\pm$ 2.7 -9.4 $\pm$ 9.2 -11.2 $\pm$ 8.8
Large SV (n=1n{=}1) 38.6 $\pm$ 14.5 2.8 $\pm$ 11.1 -10.6 $\pm$ 2.6 -4.1 $\pm$ 9.9 -0.0 $\pm$ 9.1
SV (n=5n{=}5) 50.1 $\pm$ 13.7 10.7 $\pm$ 11.8 -19.7 $\pm$ 2.3 -8.7 $\pm$ 9.1 -3.1 $\pm$ 9.0
STP (n=1n{=}1) 118.3 $\pm$ 15.8 38.1 $\pm$ 13.7 -1.2 $\pm$ 2.9 5.6 $\pm$ 9.7 -3.7 $\pm$ 9.2
MC-STP (n=5n{=}5) 97.7 $\pm$ 14.4 15.4 $\pm$ 11.7 -11.7 $\pm$ 2.5 -6.8 $\pm$ 9.2 2.6 $\pm$ 9.0
[…]
C.2.2 Uncalibrated
[…]
ARC2 Easy ARC2 Challenge MMLU MedQA MedMCQA
Base SV (n=1n{=}1) 6.5 $\pm$ 13.1 2.6 $\pm$ 13.8 -9.6 $\pm$ 2.6 -38.9 $\pm$ 98.5 -4.2 $\pm$ 12.0
SV (n=5n{=}5) 16.1 $\pm$ 21.3 -11.3 $\pm$ 18.4 -11.9 $\pm$ 2.6 -55.7 $\pm$ 82.4 -14.4 $\pm$ 18.6
STP (n=1n{=}1) 6.4 $\pm$ 344.3 -30.0 $\pm$ 211.7 -6.8 $\pm$ 2.8 -13.9 $\pm$ 87.0 -21.2 $\pm$ 159.6
MC-STP (n=5n{=}5) -42.5 $\pm$ 116.2 -39.0 $\pm$ 113.6 -9.0 $\pm$ 2.7 -19.8 $\pm$ 107.9 -42.9 $\pm$ 68.9
Large SV (n=1n{=}1) 167.9 $\pm$ 38.0 89.3 $\pm$ 36.6 -7.9 $\pm$ 2.7 9.9 $\pm$ 10.2 73.3 $\pm$ 32.8
SV (n=5n{=}5) 143.4 $\pm$ 33.6 83.1 $\pm$ 31.6 -9.1 $\pm$ 2.6 10.1 $\pm$ 10.2 63.5 $\pm$ 32.0
STP (n=1n{=}1) 129.4 $\pm$ 19.1 59.1 $\pm$ 16.5 1.6 $\pm$ 3.0 7.3 $\pm$ 9.9 30.0 $\pm$ 11.8
MC-STP (n=5n{=}5) 71.0 $\pm$ 17.2 24.0 $\pm$ 16.5 0.3 $\pm$ 2.9 4.2 $\pm$ 10.0 9.2 $\pm$ 13.2
[…]
C.2.3 Online Learning
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MMLU
(d) MedQA
(e) MedMCQA
[…]
C.3 LLama 3B $\rightarrow$ 8B
Due to the large size of the MMLU dataset and the costs associated with making predictions on it, we omit this for the (Llama3.2-3B $\rightarrow$ Llama3.1-8B) combination in this subsection.
ARC2 Easy ARC2 Challenge MedQA MedMCQA
uncal. cal. uncal. cal. uncal. cal. uncal. cal.
Base SV (n=1n{=}1) 267.7 $\pm$ 364.0 3.8 $\pm$ 96.2 322.3 $\pm$ 335.2 -11.5 $\pm$ 81.4 15.7 $\pm$ 205.7 -48.6 $\pm$ 46.9 93.6 $\pm$ 154.5 -27.8 $\pm$ 20.9
SV (n=5n{=}5) 247.2 $\pm$ 359.8 32.1 $\pm$ 100.9 352.2 $\pm$ 325.3 0.0 $\pm$ 84.4 12.5 $\pm$ 200.0 -28.2 $\pm$ 49.4 96.1 $\pm$ 146.7 -39.5 $\pm$ 20.4
STP (n=1n{=}1) 272.7 $\pm$ 338.6 34.9 $\pm$ 99.6 279.7 $\pm$ 337.5 -40.7 $\pm$ 80.9 -35.6 $\pm$ 199.1 -8.2 $\pm$ 49.3 91.7 $\pm$ 153.0 -19.2 $\pm$ 21.9
MC-STP (n=5n{=}5) 272.4 $\pm$ 325.0 44.0 $\pm$ 96.7 279.9 $\pm$ 312.6 33.0 $\pm$ 100.0 -9.2 $\pm$ 204.9 -23.2 $\pm$ 52.8 119.0 $\pm$ 154.3 -45.0 $\pm$ 20.0
Large SV (n=1n{=}1) 312.2 $\pm$ 321.9 84.2 $\pm$ 110.6 237.4 $\pm$ 277.6 -20.5 $\pm$ 84.8 -12.4 $\pm$ 136.7 -25.3 $\pm$ 49.1 81.0 $\pm$ 100.1 -23.3 $\pm$ 21.5
SV (n=5n{=}5) 257.4 $\pm$ 324.7 63.2 $\pm$ 106.4 281.6 $\pm$ 292.7 37.4 $\pm$ 83.1 4.7 $\pm$ 131.5 -32.4 $\pm$ 48.8 84.8 $\pm$ 107.0 -14.2 $\pm$ 22.1
STP (n=1n{=}1) 331.9 $\pm$ 287.7 77.0 $\pm$ 112.2 254.6 $\pm$ 291.8 -0.7 $\pm$ 83.7 -8.8 $\pm$ 126.8 -34.4 $\pm$ 46.2 95.7 $\pm$ 99.5 -38.6 $\pm$ 20.3
MC-STP (n=5n{=}5) 422.2 $\pm$ 328.8 78.2 $\pm$ 109.8 325.5 $\pm$ 306.1 -6.6 $\pm$ 78.7 65.2 $\pm$ 150.7 -52.8 $\pm$ 48.1 91.4 $\pm$ 100.3 -11.5 $\pm$ 22.1
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
C.4 Qwen 3B $\rightarrow$ 7B
Due to the large size of the MMLU dataset and the costs associated with making predictions on it, we omit this for the (Qwen-2.5-3B $\rightarrow$ Qwen-2.5-7B) combination in this subsection.
ARC2 Easy ARC2 Challenge MedQA MedMCQA
uncal. cal. uncal. cal. uncal. cal. uncal. cal.
Base SV (n=1n{=}1) 72.6 $\pm$ 67.1 80.6 $\pm$ 98.5 -35.1 $\pm$ 45.8 -3.9 $\pm$ 62.5 -37.5 $\pm$ 27.8 -46.2 $\pm$ 15.7 -21.0 $\pm$ 21.1 -40.3 $\pm$ 14.4
SV (n=5n{=}5) 3.7 $\pm$ 48.0 13.7 $\pm$ 76.4 -17.8 $\pm$ 39.0 -59.1 $\pm$ 50.3 -51.3 $\pm$ 20.8 -54.6 $\pm$ 13.2 -39.9 $\pm$ 19.0 -46.2 $\pm$ 13.5
STP (n=1n{=}1) -19.4 $\pm$ 23.5 44.7 $\pm$ 97.3 -42.1 $\pm$ 16.8 -53.6 $\pm$ 57.0 -41.5 $\pm$ 12.1 -52.0 $\pm$ 15.5 -37.5 $\pm$ 10.6 -31.5 $\pm$ 15.3
MC-STP (n=5n{=}5) -32.1 $\pm$ 21.8 -7.5 $\pm$ 75.5 -43.3 $\pm$ 18.1 -53.2 $\pm$ 53.8 -47.8 $\pm$ 11.9 -50.4 $\pm$ 13.7 -39.9 $\pm$ 11.0 -45.5 $\pm$ 14.0
Large SV (n=1n{=}1) 246.6 $\pm$ 106.2 278.9 $\pm$ 125.3 83.6 $\pm$ 56.8 43.7 $\pm$ 65.6 -9.8 $\pm$ 20.5 -46.3 $\pm$ 15.4 10.0 $\pm$ 21.6 -29.6 $\pm$ 15.6
SV (n=5n{=}5) 99.0 $\pm$ 60.4 139.1 $\pm$ 84.2 21.5 $\pm$ 38.7 27.4 $\pm$ 58.1 -34.9 $\pm$ 14.2 -43.0 $\pm$ 14.2 -13.4 $\pm$ 17.7 -30.3 $\pm$ 14.9
[…]
MC-STP (n=5n{=}5) 21.0 $\pm$ 35.6 97.2 $\pm$ 89.8 -19.4 $\pm$ 24.9 16.2 $\pm$ 56.8 -52.8 $\pm$ 11.1 -44.7 $\pm$ 14.2 -37.4 $\pm$ 12.9 -29.5 $\pm$ 14.9
[…]
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
[…]
C.5 Ablation of Different Calibration Size
We conducted an ablation study to investigate the effect of calibration size on deferral probability verification, using STP ($n=1$). The results can be found here in Table 7.
Generally, it appears that the calibration size has little influence on these magnitudes. If we examine the standard error across the various sizes, none of them is significantly better than the others. The only thing that we noted was that with a too small calibration set, we would have more diverging chains in the Bayesian Logistic Regression sampling.
Model Dataset Cal. Size = 50 Cal. Size = 100 Cal. Size = 200 Cal. Size = 500
[…]
ARC2 Challenge $87.8\pm 24.3$ $\textbf{89.3}\pm 26.2$ $68.7\pm 25.4$ $62.1\pm 28.7$
MedQA $3.0\pm 15.4$ $\textbf{1.7}\pm 16.3$ $7.7\pm 16.6$ $16.6\pm 18.6$
MedMCQA $7.5\pm 9.9$ $10.0\pm 10.1$ $8.0\pm 9.9 -$$\textbf{-5.7}\pm 10.4$
[…]
ARC2 Challenge $\textbf{45.7}\pm 13.1$ $38.1\pm 13.7$ $40.3\pm 12.8$ $43.4\pm 15.6$
MedQA $\textbf{10.1}\pm 9.7$ $5.6\pm 9.7$ $5.3\pm 10.0$ $9.8\pm 11.9$
MedMCQA $\textbf{7.4}\pm 9.5$ $3.7\pm 9.2$ $6.7\pm 9.1$ $7.0\pm 9.3$
[…]
C.6 MMLU Subject
[…]
International Law 58.66 $\pm$ 89.93
US Foreign Policy 54.20 $\pm$ 147.34
Jurisprudence 45.95 $\pm$ 76.81
Business Ethics 33.53 $\pm$ 77.09
Sociology 33.10 $\pm$ 66.72
High School Psychology 31.38 $\pm$ 37.80
High School Government And Politics 29.64 $\pm$ 35.23
Logical Fallacies 28.68 $\pm$ 87.21
World Religions 23.14 $\pm$ 68.51
Human Aging 20.88 $\pm$ 47.71
Philosophy 18.97 $\pm$ 44.90
Computer Security 16.84 $\pm$ 440.64
Miscellaneous 16.78 $\pm$ 20.86
Management 16.30 $\pm$ 61.84
High School Microeconomics 15.60 $\pm$ 27.56
High School Geography 13.66 $\pm$ 29.49
Marketing 12.23 $\pm$ 61.58
Prehistory 11.94 $\pm$ 33.99
High School Biology 11.23 $\pm$ 42.87
Security Studies 10.47 $\pm$ 50.74
Medical Genetics 8.21 $\pm$ 34.18
College Biology 8.20 $\pm$ 39.82
Professional Psychology 4.43 $\pm$ 22.52
High School US History 4.28 $\pm$ 30.17
Clinical Knowledge 3.93 $\pm$ 48.41
Formal Logic 3.25 $\pm$ 46.67
Human Sexuality 3.07 $\pm$ 65.82
All Subjects (Average) 2.52 $\pm$ 5.23
Anatomy 2.32 $\pm$ 54.35
Public Relations 1.82 $\pm$ 152.51
College Medicine 1.19 $\pm$ 46.04
Global Facts 0.70 $\pm$ 51.04
High School Macroeconomics 0.52 $\pm$ 26.42
High School European History 0.24 $\pm$ 65.82
Abstract Algebra
Nutrition -1.55 $\pm$ 28.62
Professional Accounting -1.67 $\pm$ 31.01
High School Chemistry -1.85 $\pm$ 26.87
High School Mathematics -2.46 $\pm$ 29.03
Machine Learning -2.84 $\pm$ 40.49
Moral Disputes -3.36 $\pm$ 36.88
Elementary Mathematics -3.60 $\pm$ 23.81
Conceptual Physics -3.78 $\pm$ 29.27
High School Computer Science -4.81 $\pm$ 44.25
Professional Law -6.02 $\pm$ 21.39
High School Physics -6.83 $\pm$ 21.16
College Physics -8.02 $\pm$ 28.96
Astronomy -8.60 $\pm$ 30.55
High School Statistics -10.18 $\pm$ 26.16
Econometrics -10.50 $\pm$ 36.09
Electrical Engineering -10.79 $\pm$ 41.71
College Mathematics -12.30 $\pm$ 62.39
High School World History -16.96 $\pm$ 53.38
Professional Medicine -18.75 $\pm$ 21.12
Moral Scenarios -22.65 $\pm$ 16.20
College Chemistry -22.82 $\pm$ 41.65
College Computer Science -25.63 $\pm$ 35.75
Virology $\infty$ $\pm$ $\infty$
[…]
World Religions 35.95 $\pm$ 35.25
Anatomy 34.10 $\pm$ 50.56
Virology 29.82 $\pm$ 77.02
Miscellaneous 28.80 $\pm$ 15.64
High School Psychology 15.37 $\pm$ 14.03
Clinical Knowledge 14.80 $\pm$ 23.85
Prehistory 14.27 $\pm$ 20.97
Marketing 13.94 $\pm$ 27.15
US Foreign Policy 13.70 $\pm$ 34.01
Conceptual Physics 13.37 $\pm$ 22.75
High School Geography 12.50 $\pm$ 29.38
Jurisprudence 12.22 $\pm$ 42.25
Logical Fallacies 11.77 $\pm$ 32.46
Moral Disputes 11.12 $\pm$ 21.71
High School Biology 7.66 $\pm$ 16.00
Management 6.73 $\pm$ 35.36
College Medicine 5.84 $\pm$ 26.08
International Law 4.88 $\pm$ 24.42
High School Microeconomics 4.38 $\pm$ 15.48
Human Aging 4.16 $\pm$ 39.64
High School Macroeconomics 2.51 $\pm$ 14.27
Sociology 2.05 $\pm$ 22.75
Astronomy 1.97 $\pm$ 19.43
Philosophy 0.88 $\pm$ 20.96
Electrical Engineering 0.33 $\pm$ 26.12
Nutrition 0.31 $\pm$ 19.88
College Biology 0.30 $\pm$ 23.37
Abstract Algebra
Computer Security -0.26 $\pm$ 31.92
All Subjects (Average) -1.16 $\pm$ 2.87
Medical Genetics -1.39 $\pm$ 27.97
Business Ethics -1.59 $\pm$ 35.09
High School Government And Politics -1.61 $\pm$ 16.63
Formal Logic -2.01 $\pm$ 43.78
Professional Psychology -2.29 $\pm$ 13.79
High School Computer Science -2.58 $\pm$ 29.56
Human Sexuality -3.19 $\pm$ 19.56
Security Studies -4.89 $\pm$ 31.41
Public Relations -5.57 $\pm$ 61.18
Econometrics -5.67 $\pm$ 29.45
College Computer Science -7.06 $\pm$ 29.54
High School Statistics -7.70 $\pm$ 15.34
High School Physics -9.27 $\pm$ 25.92
College Physics -9.43 $\pm$ 15.95
Moral Scenarios -9.63 $\pm$ 13.23
Professional Medicine -10.05 $\pm$ 18.46
Elementary Mathematics -10.43 $\pm$ 8.50
Professional Accounting -10.69 $\pm$ 17.82
High School World History -10.72 $\pm$ 19.96
High School US History -11.81 $\pm$ 17.55
High School Mathematics -12.64 $\pm$ 11.41
High School Chemistry -13.66 $\pm$ 17.31
Professional Law -14.33 $\pm$ 10.24
High School European History -17.91 $\pm$ 24.25
College Chemistry -17.92 $\pm$ 38.05
College Mathematics -17.95 $\pm$ 26.48
Machine Learning -20.64 $\pm$ 19.84
Global Facts -35.32 $\pm$ 36.23
[…]
C.7 Grid-Search over Threshold Parameters
we perform an additional experiment on the ARC2-Easy dataset with the (Qwen-2.5-1.5B $\rightarrow$ Qwen-2.5-7B) combination. We perform a grid search with every parameter $\mathbf{\theta}=\{\phi_{\text{base}},\xi_{\text{base}},\xi_{\text{large}}\}$ over $\{0.5,0.15,0.25,0.35,0.45,0.55,0.65,0.75,0.85,0.95\}$. The search grid’s time complexity is cubic, $O(n^{3})$, and increases with the addition of expert data to the replay buffer, becoming computationally extremely expensive compared to the gradient-based approach. We report our findings in Figure 16. We observe from the results, that the gradient-based approach achieves lower cumulative regret, compared to the grid-search approach, be it on single-model strategies, and also on the cascsaded LLM framework.
(a) Gradient-based Optimisation
(b) Grid Search
Figure 16: Gradient-Based vs. Grid-Search. Online learning performance of the cascaded LLM framework over 1000 samples using the proposed gradient-based (16(a)) approach against a grid-search (16(b)) over the threshold parameters $\mathbf{\theta}$.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: For assumptions see section 3, for results see section 5, where the experiments represent the claims made in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of our methods and experiments in section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: In this paper there are no theoretical results. It is purely empirical.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide the full set of description of the experiments in section 4 and section 5. Moreover, we provide our code: https://anonymous.4open.science/r/helm-82D7
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide our code: https://anonymous.4open.science/r/helm-82D7 and all the models and datasets in our experiments are open-source and available for everyone.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: To the best of our knowledge, weWe provide the full set of hyperparameters in section 5. Moreover, we provide our code with all the experimental config files: https://github.com/fanconic/cascaded-llms
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In section 5 we report all the tables and figures with the standard error and 95%-confidence intervals.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, in section 5 we point out the available access to an A100 GPU.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: To the best of my knowledge, we do this.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper presents work whose goal is to advance the field of Machine Learning.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release any data or model with this paper
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: To the best of our knowledge we have credited the original owners
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: no research with crowdsourcing or human subjects
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: no research with crowdsourcing or human subjects
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: the core method development in this research does not involve LLMs as any important, original, or non-standard components
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.